### v1

# The general Brannan coefficient conjecture and Watson’s lemma

Abstract
The coefficients $A_{n}(\alpha,\beta,\omega)$ in the Maclaurin expansion $(1+\omega z)^{\alpha}(1-z)^{-\beta}=\sum_{n=0}^{\infty}A_{n}(\alpha,\beta,\omega)z^{n}$ are studied, where $\omega,z\in\mathbb{C}$ with $|z|<|\omega|=1$, and $\alpha,\beta\in(0,1]$. In 1973 Brannan conjectured that $|A_{n}(\alpha,\beta,\omega)|\leq A_{n}(\alpha,\beta,1)$ for each positive odd integer $n$, and showed it is true for $n=3$. This has recently been proven for all odd integers $n\geq 5$ by a number of authors in aggregate for the special case $\beta=1$. In this paper hypergeometric integral representations and Watson-type approximations are utilised, from which the general problem is reduced to numerically evaluating the minima of certain simple, explicit, slowly-varying functions over compact domains. From the positivity of these constants it is shown that the conjecture holds for $\alpha,\beta\in(0,1]$, $0\leq|\arg(\omega)|\leq\pi-\phi_{0}$ and $n=5,7,9,\ldots$, where $\phi_{0}=0.061$.
keywords:
Brannan’s conjecture, Watson’s lemma, hypergeometric functions, univalent functions, coefficient inequalities
{AMS}
33C05, 30C45, 30C50, 41A60
1 Introduction
Consider the Maclaurin series of the following univalent function
(1) $\frac{(1+\omega z)^{\alpha}}{(1-z)^{\beta}}=\sum_{n=0}^{\infty}A_{n}(\alpha,\beta,\omega)z^{n},$
where $z,\omega\in\mathbb{C}$ with $|z|<1=|\omega|$, and $\alpha,\beta\in(0,1]$. Each $A_{n}(\alpha,\beta,\omega)$ is a polynomial in $\omega$ of degree $n$, and it is well known that
(2) $A_{n}(\alpha,\beta,\omega)=\frac{(\beta)_{n}}{n!}\,{}_{2}F_{1}(-n,-\alpha;\,1-\beta-n;\,-\omega),$
where $_{2}F_{1}$ is the Gauss hypergeometric function given by
(3) $_{2}F_{1}(a,b;\,c;\,z)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{(c)_{k}\,k!}\,z^{k},$
for $|z|<1$, and elsewhere by analytic continuation, with the Pochhammer symbol defined in the usual way
(4) $(a)_{n}=\frac{\Gamma(a+n)}{\Gamma(a)}.$
The following conjecture due to Brannan [Brannan:1974:OCP] in 1973 arose in relation to coefficient problems involving functions of bounded boundary rotation.
Brannan’s Conjecture. Suppose $\alpha,\beta\in(0,1]$ and $0\leq|\theta|\leq\pi$. Then
(5) $\left|A_{n}(\alpha,\beta,e^{i\theta})\right|\leq A_{n}(\alpha,\beta,1)\quad(n=3,5,7,\ldots).$
In this paper we prove this to be true for $0\leq|\theta|\leq\pi-0.061$ and $n=5,7,9,\ldots$; Brannan [Brannan:1974:OCP] had already shown it to be true for $n=3$.
Remark 1.
From [Barnard:2015:BCA] it was shown for $\alpha,\beta\in(0,1)$ and positive integer $n$
(6) $A_{n}(\alpha,\beta,\omega)=\frac{\Gamma(n+\beta)}{\Gamma(\beta)\,n!}+\frac{\Gamma(n-\alpha+\beta)\Gamma(1+\alpha)\sin(\pi\alpha)}{\pi\Gamma(\beta)\,n!}\\ \times\int_{0}^{1}p_{n}(\omega,t)\,t^{-1-\alpha}(1-t)^{n-1+\beta}\,dt,$
where
(7) $p_{n}(\omega,t)=1-\left(1-\frac{\omega t}{1-t}\right)^{n},$
and from this it is straightforward to show that $A_{n}(\alpha,\beta,1)>0$ if $n$ is odd; the same is not necessarily true if $n$ is even.
As we mentioned, Brannan [Brannan:1974:OCP] verified his conjecture for the first nontrivial odd case $n=3$, and also showed that the inequality does not hold in general for even $n$. Aharonov and Friedland [Aharonov:1972:OAI] proved the inequality Eq. 5 in the case $\alpha,\beta\geq 1$.
For the special case $\beta=1$ and $\alpha\in(0,1)$, early proofs were obtained for the individual odd indices $n=5$ (Milcetich [Milcetich:1989:OAC]), and $n=7$ (Barnard, Pearce and Wheeler [Barnard:1997:OAC]). Jayatilake [Jayatilake:2013:BCF] later verified the conjecture for all odd $n\leq 51$ using a computer-assisted squaring procedure.
More recently, Szász [Szasz:2020:OTB] and (independently) Deniz, Çaglar and Szász [Deniz:2020:TFS] proved complementary large-$n$ ranges via integral representations, giving a cumulative proof of Brannan’s conjecture for $\beta=1$. Barnard and Richards [Barnard:2021:ADP] subsequently provided a fully direct analytic proof for $\beta=1$ valid for all positive odd $n$.
In the general two-parameter setting, Ruscheweyh and Salinas [Ruscheweyh:2007:OBC] established the conjecture for all odd $n$ on the diagonal $\alpha=\beta\in(0,1)$. Recently, Cotîrla and Szász [Cotirla:2024:OTG] in 2024 proved the conjecture in the region $3/4\leq\alpha\leq\beta\leq 1$ for all odd indices, using a new integral representation.
The plan of this paper is as follows. In Section 2 we first express $A_{n}(\alpha,\beta,\omega)$ in an alternative hypergeometric form, from which we derive a loop-integral representation. This is then decomposed into a sum of two Laplace-type integrals. These are employed to obtain a quantitative lower bound for the difference between the left- and right-hand sides of Eq. 5. The resulting representation yields a tractable lower bound for the difference at $\omega=1$ and, after taking moduli, for $\omega$ on the unit circle.
We then in Section 3 apply the method of Watson’s lemma (see, for example, [Olver:1997:ASF, Chap. 3], [Temme:2015:AMF, Chap. 2], [Wong:1989:AAI, Chap. 5]) to derive a two-term large-$n$ asymptotic approximation for the lower bound, together with explicit remainders. These remainder terms are given by integral representations, and from these we obtain simple lower bounds by computing the infima of certain smooth, explicit, slowly-varying functions over compact sets.
In Section 4 we prove the conjecture for $n=5$, with $\alpha,\beta\in(0,1]$ and $0\leq\arg(\omega)\leq\pi-0.061$, by establishing positivity of the lower bound constructed from the main terms and the explicit lower error bounds of Section 3.
In Section 5 we extend the result to all odd $n\geq 7$. This is done by proving positivity of an appropriate derivative, again reducing the required estimates to minimizing a pair of simple, explicit, slowly-varying functions over a compact domain. Finally, in Section 6 we discuss a possible way of tackling the missing segment $\pi-0.061<|\arg(\omega)|\leq\pi$.
We emphasise that all the functions being minimized, with the exception of the real-valued gamma function $\Gamma(x)$ with $x\in(0,2]$, and the real-valued incomplete gamma function $\Gamma(x,5)$ with $x\in[3,4]$ (both of which are readily computable to a high degree of accuracy), only involve slowly-varying elementary functions, and do not involve large/infinite sums or integrals. The only minor complications are mild removable singularities at the boundaries of the regions under consideration, and these are easily handled numerically by a limiting process.
2 Integral lower bounds
From Euler’s reflection formula
(8) $\frac{1}{\Gamma(z)\Gamma(1-z)}=\frac{\sin(\pi z)}{\pi},$
Eq. 2 and [NIST:DLMF, Eqs. 5.2.5, 5.2.6, 15.8.6] we obtain for positive integer $n$
(9) $A_{n}(\alpha,\beta,\omega)=(-1)^{n+1}\frac{\Gamma(1+\alpha)\Gamma(n-\alpha)\sin(\pi\alpha)}{\pi n!}\\ \times\omega^{n}\ {}_{2}F_{1}\left(\beta,-n;\alpha-n+1;-1/\omega\right).$
Letting $x=-1/\omega=\exp(i\phi)$, so that
(10) $\phi=\pi-\theta\quad(\mathrm{mod}\;2\pi),$
and then appealing to the Schwarz reflection principle, we see that Brannan’s conjecture is equivalent to
(11) $w_{n}(\alpha,\beta,-1)\geq\left|w_{n}(\alpha,\beta,e^{i\phi})\right|,$
for $\alpha,\beta\in(0,1]$, $\phi\in[0,\pi]$ and $n=3,5,7,\ldots$, where
(12) $w_{n}(\alpha,\beta,x)=\sin(\pi\alpha)\,{}_{2}F_{1}\!\left(\beta,-n;\alpha-n+1;x\right).$
Note from Eqs. 1, 4, 9, and 12 that for $n=1,2,3,\ldots$
(13) $\lim_{\alpha\to 0}w_{n}(\alpha,\beta,x)=-\frac{\pi\Gamma(\beta+n)}{\Gamma(\beta)(n-1)!}x^{n},$
and for $n=2,3,4\ldots$
(14) $\lim_{\alpha\to 1}w_{n}(\alpha,\beta,x)=\frac{\pi\,\Gamma(\beta+n-1)}{\Gamma(\beta)(n-2)!}\left\{n-(\beta+n-1)x\right\}\,x^{n-1}.$
We wish to attack the problem using Watson’s lemma for large $n$. The following representation will enable us to do so. In this, and throughout, all complex powers use the principal branches of logarithms.
Lemma 2.1.
Assume $\alpha,\beta\in[0,1]$, $1\leq n\in\mathbb{N}$, and $x\in\mathbb{C}\setminus[0,\infty)$. Then
(15) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)}{n!}w_{n}(\alpha,\beta,x)\\ =(-1)^{n+1}\sin(\pi\beta)(-x)^{\,n-\alpha}\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,(e^{s}-x)^{\alpha}e^{-ns}\,ds\\ +\sin(\pi\alpha)(-x)^{-\beta}\int_{0}^{\infty}(e^{s}-1)^{\alpha}\left(e^{s}-1/x\right)^{-\beta}e^{-ns}ds.$
When $\beta=1$ it is understood that the limit of the first term on the RHS applies.
Proof 2.2.
From [NIST:DLMF, Eqs. 15.1.2 and 15.6.4]
(16) $_{2}F_{1}(a,b;c;z)=\frac{e^{-b\pi i}\Gamma(1-b)\Gamma(c)}{2\pi i\,\Gamma(c-b)}\int_{1}^{(0+)}\frac{t^{\,b-1}(1-t)^{c-b-1}}{(1-zt)^{a}}\,dt,$
with $b\neq 1,2,3,\dots$, $c\neq 0,-1,-2,\dots$, $\Re(c-b)>0$, and $|\arg(1-z)|<\pi$. Here the contour is the standard loop starting at $t=1$ winding once counterclockwise around $t=0$ and returning to $t=1$.
So with $a=\beta$, b=nb=-n, $c=1+\alpha-n$ and $x=1/\xi$ we have from Eqs. 8 and 12 the integral representation
(17) $w_{n}(\alpha,\beta,1/\xi)=\frac{i\,n!}{2\Gamma(n-\alpha)\Gamma(1+\alpha)}(-\xi)^{\beta}I(\xi),$
where
(18) $I(\xi)=\int_{1}^{(0+)}\frac{t^{-n-1}(1-t)^{\alpha}}{(t-\xi)^{\beta}}\,dt.$
We temporarily assume $\xi=e^{i\theta}$ with $\theta\in(0,\pi]$ and $\alpha,\beta\in(0,1)$, and for these values the integral converges, $\xi$ lies off the cut $1,\infty)$, and the branch points $t=1$ and $t=\xi$ are distinct. The contour initially lies inside the unit circle $|t|=1$ and so avoids the branch point $t=\xi$. We take the branch cut for $(1-t)^{\alpha}$ to be $1,\infty)$ (principal branch on $\mathbb{C}\setminus[1,\infty)$), and the branch cut for $(t-\xi)^{-\beta}$ to be the ray extending from $\xi$ to $\infty$ with $\arg(t)=\arg(\xi)=\theta$.
Since $\alpha,\beta\in(0,1)$ the branch points at $t=1$ and $t=\xi$ are locally integrable, and as $t\to\infty$ the integrand is $\mathcal{O}(t^{-n-1+\alpha-\beta})$ so contributions from circular arcs vanish at infinity. Hence the loop in Eq. 18 may be deformed to four rays, two on either side of both of the cuts. As a result we arrive at
(19) $I(\xi)=I_{\xi}^{+}(\xi)-I_{\xi}^{-}(\xi)+I_{1}^{+}(\xi)-I_{1}^{-}(\xi),$
where $I_{\xi}^{\pm}(\xi)$ denotes the integral of $t^{-n-1}(1-t)^{\alpha}(t-\xi)^{-\beta}$ from $t=\xi$ to $t=\infty\exp(i\theta)$ with $\arg(\xi-t)=\arg(\xi)\mp\pi$, and $I_{1}^{\pm}(\xi)$ denotes the integral of the same integrand from $t=1$ to $t=\infty$ with $\arg(1-t)=\mp\pi$.
For the first pair make the change of variable $t=\xi v$ with $1\leq v<\infty$, so that for $I_{\xi}^{\pm}(\xi)$
$(t-\xi)^{-\beta}=(-\xi)^{-\beta}(1-v)^{-\beta}=e^{\pm i\pi\beta}(-\xi)^{-\beta}(v-1)^{-\beta}\quad(1\leq v<\infty).$
Therefore, since $(1-t)^{\alpha}=(1-\xi v)^{\alpha}$ takes the same value for both integrals for $\theta\in(0,\pi)$, we obtain
(20) $I_{\xi}^{+}(\xi)-I_{\xi}^{-}(\xi)=2i(-1)^{n}\sin(\pi\beta)(-\xi)^{-n-\beta}\int_{1}^{\infty}v^{-n-1}(v-1)^{-\beta}(1-\xi v)^{\alpha}\,dv.$
Now for $v$ above and below the cut $1,\infty)$
$(1-v)^{\alpha}=e^{\mp i\pi\alpha}(v-1)^{\alpha},$
and therefore (with $t$ replaced by $v$ for consistency)
(21) $I_{1}^{+}(\xi)-I_{1}^{-}(\xi)=-2i\sin(\pi\alpha)\int_{1}^{\infty}v^{-n-1}(v-1)^{\alpha}(v-\xi)^{-\beta}dv.$
Inserting Eqs. 20 and 21 into Eq. 19 yields
(22) $I(\xi)=2i(-1)^{n}\sin(\pi\beta)(-\xi)^{-n-\beta}\int_{1}^{\infty}v^{-n-1}(v-1)^{-\beta}(1-\xi v)^{\alpha}\,dv\\ -2i\sin(\pi\alpha)\int_{1}^{\infty}v^{-n-1}(v-1)^{\alpha}(v-\xi)^{-\beta}dv.$
Now plug Eq. 22 into Eq. 17 and in each of the resulting integrals set $v=e^{s}$, so that $dv=e^{s}\,ds$ and $v^{-n-1}dv=e^{-ns}\,ds$. Then, on using $(1-\xi e^{s})^{\alpha}=(-\xi)^{\alpha}(e^{s}-\xi^{-1})^{\alpha}$ and replacing $\xi$ by $1/x$, we confirm that Eq. 15 holds for $\alpha,\beta\in(0,1)$ and $\arg(1/x)=\theta\in(0,\pi]$. Finally, the identity extends to $\alpha,\beta\in[0,1]$ by continuity (all boundary singularities are removable, with the case $\beta=1$ interpreted by the limiting value of the first term), and it extends from $x=e^{-i\theta}$, $\theta\in(0,\pi]$, to all $x\in\mathbb{C}\setminus[0,\infty)$ by analytic continuation in $x$.
Our next step is to use this representation to give a suitable lower bound for the difference of the two terms in Eq. 11.
Lemma 2.3.
Let $\alpha,\beta\in(0,1]$, $1\leq n\in\mathbb{N}$ be odd, $\phi\in[0,\pi)$, and define
(23) $\Delta=\pi-\phi>0,$
(24) $R(s,\phi)=\left|e^{s}-e^{\pm i\phi}\right|=\sqrt{e^{2s}+1-2e^{s}\cos(\phi)},$
(25) $I_{1}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{-\beta}\left\{\left(e^{s}+1\right)^{\alpha}-R(s,\phi)^{\alpha}\right\}e^{-ns}ds,$
(26) $I_{2}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{\alpha}\left\{\left(1+e^{s}\right)^{-\beta}-R(s,\phi)^{-\beta}\right\}e^{-ns}ds,$
and
(27) $h(\alpha,\beta,\phi;n)=\sin(\pi\beta)\,I_{1}(\alpha,\beta,\phi;n)+\sin(\pi\alpha)\,I_{2}(\alpha,\beta,\phi;n).$
Then
(28) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}}{n!\Delta^{2}}\,\geq h(\alpha,\beta,\phi;n).$
Remark 2.4.
We introduced the factor $\Delta^{-2}$ for convenience, in order that both sides of the inequality Eq. 28 do not vanish as $\phi\to\pi$, while remaining bounded in this limit, as we show shortly.
Proof 2.5.
In Eq. 15 set x=1x=-1 and $x=\exp(i\phi)$, taking absolute values and then the integral triangle inequality on the latter. On subtraction of the two we deduce that
(29) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)}{n!}\,\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}\\ \geq\sin(\pi\beta)\,\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,\left\{(e^{s}+1)^{\alpha}-\left|e^{s}-e^{i\phi}\right|^{\alpha}\right\}e^{-ns}\,ds\\ \quad+\sin(\pi\alpha)\,\int_{0}^{\infty}(e^{s}-1)^{\alpha}\,\left\{(e^{s}+1)^{-\beta}-\left|e^{s}-e^{-i\phi}\right|^{-\beta}\right\}e^{-ns}\,ds,$
and Eq. 28 then follows on division of both sides by $\Delta^{2}$.
3 Two-term Watson-type expansions for $I_{1}$ and $I_{2}$
Our Watson-type approximation with explicit remainder terms reads as follows.
Lemma 3.1.
Let
(30) $C_{1}(\alpha,\phi)=\frac{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}{\Delta^{2}},$
(31) $C_{2}(\beta,\phi)=\frac{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}{\Delta^{2}},$
(32) $K_{1}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{s}{e^{s}-1}\right)^{\beta}\frac{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}}{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}-1-\frac{\alpha-\beta}{2}\,s\right],$
and
(33) $K_{2}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{e^{s}-1}{s}\right)^{\alpha}\frac{R(s,\phi)^{-\beta}-(1+e^{s})^{-\beta}}{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}-1-\frac{\alpha-\beta}{2}\,s\right],$
with $R(s,\phi)$ defined by Eq. 24. Then, for $\alpha,\beta\in(0,1)$ and $\phi\in[\phi_{0},\pi)$ for fixed $\phi_{0}>0$,
(34) $I_{1}(\alpha,\beta,\phi;n)=\frac{C_{1}(\alpha,\phi)\,\Gamma(1-\beta)}{n^{1-\beta}}\left\{1+\frac{(\alpha-\beta)(1-\beta)}{2n}\right\}+E_{1}(\alpha,\beta,\phi;n),$
and
(35) $I_{2}(\alpha,\beta,\phi;n)=-\frac{C_{2}(\beta,\phi)\,\Gamma(1+\alpha)}{n^{1+\alpha}}\left\{1+\frac{(\alpha-\beta)(1+\alpha)}{2n}\right\}+E_{2}(\alpha,\beta,\phi;n),$
where
(36) $E_{1}(\alpha,\beta,\phi;n)=C_{1}(\alpha,\phi)\int_{0}^{\infty}s^{2-\beta}K_{1}(\alpha,\beta,\phi;s)e^{-ns}\,ds,$
and
(37) $E_{2}(\alpha,\beta,\phi;n)=-C_{2}(\beta,\phi)\int_{0}^{\infty}s^{2+\alpha}K_{2}(\alpha,\beta,\phi;s)e^{-ns}\,ds.$
Remark 3.2.
We have
(38) $K_{1}(\alpha,\beta,\phi;s),\ K_{2}(\alpha,\beta,\phi;s)=\mathcal{O}(1)\quad\text{as }s\to 0,$
uniformly for $\alpha,\beta\in(0,1)$ and $\phi\in[\phi_{0},\pi)$. In particular, all small-$s$ asymptotics are uniform in $\phi$ up to $\phi\to\pi$. The following limits as $\phi\to\pi$ and $\alpha+\beta\to 0$ also apply:
(39) $K_{1}(\alpha,\beta,\pi;s)=\frac{1}{s^{2}}\Biggl[\left(\frac{s}{e^{s}-1}\right)^{\beta}\frac{4e^{s}}{(e^{s}+1)^{2}}\left(\frac{e^{s}+1}{2}\right)^{\alpha}-1-\frac{\alpha-\beta}{2}\,s\Biggr],$
(40) $K_{2}(\alpha,\beta,\pi;s)=\frac{1}{s^{2}}\Biggl[\left(\frac{e^{s}-1}{s}\right)^{\alpha}\frac{4e^{s}}{(e^{s}+1)^{2}}\left(\frac{2}{e^{s}+1}\right)^{\beta}-1-\frac{\alpha-\beta}{2}\,s\Biggr],$
and
(41) $K_{1}(0,0,\phi;s)=K_{2}(0,0,\phi;s)=\frac{1}{s^{2}}\left[\frac{\ln\left\{\left(1+e^{s}\right)/R(s,\phi)\right\}}{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}-1\right]\quad(0<\phi<\pi).$
Furthermore,
(42) $K_{1}(\alpha,\beta,\phi;s)=-\frac{\alpha-\beta}{2s}-\frac{1}{s^{2}}+\mathcal{O}\left(s^{\beta-2}e^{-(1-\alpha+\beta)s}\right),$
and
(43) $K_{2}(\alpha,\beta,\phi;s)=-\frac{\alpha-\beta}{2s}-\frac{1}{s^{2}}+\mathcal{O}\left(s^{-\alpha-2}e^{-(1-\alpha+\beta)s}\right),$
as $s\to\infty$.
The coefficients $C_{1}(\alpha,\phi)$ and $C_{2}(\beta,\phi)$ defined in Eqs. 30 and 31 remain smooth and strictly positive for $\phi\in[\phi_{0},\pi)$ and $\alpha,\beta\in(0,1)$, and have finite nonzero limits as $\phi\to\pi$; in particular
(44) $C_{1}(\alpha,\phi)=\frac{2^{\alpha}\,\alpha}{8}+\mathcal{O}\left\{(\phi-\pi)^{2}\right\},$
and
(45) $C_{2}(\beta,\phi)=\frac{2^{-\beta}\,\beta}{8}+\mathcal{O}\left\{(\phi-\pi)^{2}\right\},$
as $\phi\to\pi$.
Proof 3.3.
From Eqs. 25 and 26
(46) $I_{j}(\alpha,\beta,\phi;n)=\int_{0}^{\infty}G_{j}(\alpha,\beta,\phi;s)e^{-ns}ds\quad(j=1,2),$
where
(47) $G_{1}(\alpha,\beta,\phi;s)=\frac{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}}{\Delta^{2}(e^{s}-1)^{\beta}},$
and
(48) $G_{2}(\alpha,\beta,\phi;s)=\frac{(e^{s}-1)^{\alpha}}{\Delta^{2}}\left[(1+e^{s})^{-\beta}-R(s,\phi)^{-\beta}\right].$
Using a Taylor expansion at $s=0$ and the definition Eqs. 30 and 31 one finds that, as $s\to 0$,
$s^{\beta}G_{1},\,s^{-\alpha}G_{2}=\pm C_{1,2}\left\{1+\tfrac{1}{2}(\alpha-\beta)s+\mathcal{O}\left(s^{2}\right)\right\},$
uniformly for $\alpha,\beta\in(0,1)$ and $\phi\in[\phi_{0},\pi)$. We therefore can write the integrands of Eq. 46 in the forms
(49) $G_{1}(\alpha,\beta,\phi;s)=C_{1}(\alpha,\phi)s^{-\beta}\left\{1+\tfrac{1}{2}(\alpha-\beta)s+s^{2}K_{1}(\alpha,\beta,\phi;s)\right\},$
and
(50) $G_{2}(\alpha,\beta,\phi;s)=-C_{2}(\beta,\phi)s^{\alpha}\left\{1+\tfrac{1}{2}(\alpha-\beta)s+s^{2}K_{2}(\alpha,\beta,\phi;s)\right\},$
where the remainder kernels $K_{1,2}(\alpha,\beta,\phi;s)$ are given by Eqs. 32 and 33 and satisfy Eq. 38. Substituting Eqs. 49 and 50 into Eqs. 25 and 26, and using
(51) $\int_{0}^{\infty}s^{\mu}e^{-ns}\,ds=\frac{\Gamma(1+\mu)}{n^{1+\mu}}\quad(\Re(n)>0,\Re(\mu)>-1),$
we obtain Eqs. 34 and 35.
Let us now consider the RHS of Eq. 28, suitably scaled. We introduce the prefactors
(52) $A_{1}(\alpha,\phi)=\frac{(1-\alpha)C_{1}(\alpha,\phi)}{\sin(\pi\alpha)\Gamma(1+\alpha)}=\frac{(1-\alpha)\left[2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}\right]}{\sin(\pi\alpha)\Gamma(1+\alpha)(\pi-\phi)^{2}},$
and
(53) $A_{2}(\alpha,\beta,\phi)=\frac{(1-\alpha)C_{2}(\beta,\phi)}{\sin(\pi\beta)\Gamma(1-\beta)}=\frac{(1-\alpha)\Gamma(\beta)\left[\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}\right]}{\pi(\pi-\phi)^{2}}.$
Then, on combining Eqs. 27, 34, 35, 36, 37, 52, and 53, one arrives at:
Theorem 3.4.
For $\alpha,\beta\in[0,1]$ and $\phi\in(0,\pi]$
(54) $\mathcal{H}(\alpha,\beta,\phi;n)=\mathcal{H}^{(2)}(\alpha,\beta,\phi;n)+\mathcal{E}_{0}(\alpha,\beta,\phi;n)+\mathcal{E}_{\infty}(\alpha,\beta,\phi;n),$
where
(55) $\mathcal{H}(\alpha,\beta,\phi;n)=\frac{(1-\alpha)n^{1+\alpha-\beta}h(\alpha,\beta,\phi;n)}{(\alpha+\beta)\sin(\pi\alpha)\sin(\pi\beta)\Gamma(1-\beta)\Gamma(1+\alpha)},$
(56) $\mathcal{H}^{(2)}(\alpha,\beta,\phi;n)=\frac{1}{\alpha+\beta}\left[A_{1}(\alpha,\phi)n^{\alpha}\left\{1+\frac{(\alpha-\beta)(1-\beta)}{2n}\right\}\right.\\ \left.-\frac{A_{2}(\alpha,\beta,\phi)}{n^{\beta}}\left\{1+\frac{(\alpha-\beta)(1+\alpha)}{2n}\right\}\right],$
(57) $\mathcal{E}_{0}(\alpha,\beta,\phi;n)=n^{1+\alpha-\beta}\int_{0}^{1}L(\alpha,\beta,\phi;s)e^{-ns}ds,$
and
(58) $\mathcal{E}_{\infty}(\alpha,\beta,\phi;n)=n^{1+\alpha-\beta}\int_{1}^{\infty}L(\alpha,\beta,\phi;s)e^{-ns}ds,$
in which
(59) $L(\alpha,\beta,\phi;s)=\frac{s^{2}}{\alpha+\beta}\left\{\frac{A_{1}(\alpha,\phi)K_{1}(\alpha,\beta,\phi;s)}{\Gamma(1-\beta)s^{\beta}}-\frac{s^{\alpha}A_{2}(\alpha,\beta,\phi)K_{2}(\alpha,\beta,\phi;s)}{\Gamma(1+\alpha)}\right\},$
and in all cases (finite) limits applying when $\alpha\to 0,1$, $\beta\to 0,1$ and/or $\phi\to\pi$. In Eq. 59 $K_{1}(\alpha,\beta,\phi;s)$ and $K_{2}(\alpha,\beta,\phi;s)$ are given by Eqs. 24, 32, and $3$.
Note, from Eqs. 52 and 53,
(60) $A_{1}(0,\phi)=\frac{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}{\pi(\pi-\phi)^{2}},$
(61) $A_{1}(1,\phi)=\frac{2\left\{1-\sin(\tfrac{1}{2}\phi)\right\}}{\pi\Delta^{2}},$
(62) $A_{1}(\alpha,\pi)=\frac{2^{\alpha}\Gamma(2-\alpha)}{8\pi},$
(63) $A_{2}(\alpha,0,\phi)=\frac{(1-\alpha)\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}{\pi(\pi-\phi)^{2}},$
(64) $A_{2}(\alpha,1,\phi)=\frac{(1-\alpha)\left\{\csc(\tfrac{1}{2}\phi)-1\right\}}{2\pi\Delta^{2}},$
and
(65) $A_{2}(\alpha,\beta,\pi)=\frac{(1-\alpha)\Gamma(1+\beta)}{8\pi\,2^{\beta}},$
and $L(\alpha,\beta,\pi;s)$ follows from Eqs. 39, 40, 59, 62, and 65.
Furthermore, using the $\mathcal{O}(1)$ behaviour of $K_{1}(\alpha,\beta,\phi;s)$ and $K_{2}(\alpha,\beta,\phi;s)$ as $s\to 0$, one finds that
(66) $L(\alpha,\beta,\phi;s)=\mathcal{O}\left(s^{2-\beta}\right)\quad(s\to 0),$
for $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi]$. Note from Eq. 53 and $1/\Gamma(0)=0$ that $L(1,1,\phi;s)=0$. Moreover the limit $L(0,0,\phi;s)$ exists and is finite for all $\phi\in[\phi_{0},\pi)$ and bounded $s\geq 0$. In particular, we have
(67) $L(0,0,\phi;s)=\frac{1}{2\pi(\pi-\phi)^{2}}\Biggl[2\ln\left\{\frac{R(s,\phi)}{e^{s}+1}\right\}\ln\left(e^{s}-1\right)+\ln^{2}\left(e^{s}+1\right)\biggr.\\ \left.-\ln^{2}\left\{R(s,\phi)\right\}+\ln\left\{\sin(\tfrac{1}{2}\phi)\right\}\ln\left\{\frac{4\sin(\tfrac{1}{2}\phi)}{s^{2}}\right\}\right]\\ =\frac{\left\{1+\cos(\phi)\right\}s^{2}\ln(s)}{8\pi\left\{1-\cos(\phi)\right\}(\pi-\phi)^{2}}+\mathcal{O}\left(s^{2}\right)\quad(s\to 0),$
where $R(s,\phi)$ is given by Eq. 24. From this
(68) $L(0,0,\pi;s)=\frac{1}{8\pi}\left[\mathrm{sech}^{2}(\tfrac{1}{2}s)\ln\left\{\coth(\tfrac{1}{2}s)\right\}+\ln(\tfrac{1}{2}s)\right]\\ =\frac{s^{2}\ln(s)}{32\pi}+\mathcal{O}\left(s^{2}\right)\quad(s\to 0).$
We also note from Eqs. 23, 24, 32, 33, 59, 61, and 64 that
(69) $L(0,1,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\frac{\csc(\tfrac{1}{2}\phi)-1}{2}\left(1-\frac{s}{2}\right)-\frac{1}{R(s,\phi)}+\frac{1}{1+e^{s}}\right\},$
(70) $L(1,0,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\sin(\tfrac{1}{2}\phi)(s+2)-R(s,\phi)+e^{s}-s-1\right\},$
and from Eqs. 42, 43, and 59
(71) $L(\alpha,\beta,\phi;s)\sim\frac{\alpha-\beta}{2(\alpha+\beta)}\left\{\frac{A_{2}(\alpha,\beta,\phi)}{\Gamma(1+\alpha)}\,s^{1+\alpha}-\frac{A_{1}(\alpha,\phi)}{\Gamma(1-\beta)}\,s^{1-\beta}\right\}\quad(s\to\infty).$
Our task now is to bound, from below, the remainder terms $\mathcal{E}_{0}(\alpha,\beta,\phi;n)$ and $\mathcal{E}_{\infty}(\alpha,\beta,\phi;n)$. The following constants play an important role in this.
Proposition 3.5.
For $\phi_{0}=0.061$ let
(72) $m_{0}=\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi),\ s\in[0,1]\end{subarray}}L_{0}(\alpha,\beta,\phi;s),$
where
(73) $L_{0}(\alpha,\beta,\phi;s)=s^{-1+\beta}\,L(\alpha,\beta,\phi;s),$
with $L_{0}(\alpha,\beta,\phi;0)=\lim_{s\to 0}L_{0}(\alpha,\beta,\phi;s)=0$, and
(74) $m_{\infty}=\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi),\ s\in[1,\infty)\end{subarray}}L_{\infty}(\alpha,\beta,\phi;s),$
where
(75) $L_{\infty}(\alpha,\beta,\phi;s)=s^{-2-\alpha}\,L(\alpha,\beta,\phi;s).$
Then
(76) $m_{0}=-0.0976382271\cdots,$
attained at $\alpha=0$, $\beta=0$, $\phi=\phi_{0}$, $s=s_{0}:=0.0215923350\cdots$ (cf. Eq. 67), and
(77) $m_{\infty}=-0.03332478558\cdots,$
attained at $\alpha=0$, $\beta=1$, $\phi=\phi_{0}$, $s=s_{\infty}:=3.999154909\cdots$ (cf. Eq. 69).
Remark 3.6.
On the compact set $\alpha,\beta\in[0,1]$, $\phi\in[\phi_{0},\pi]$ and $s\in[0,1]$ the function $L_{0}(\alpha,\beta,\phi;s)$ is continuous, bounded and $\mathcal{O}(s)$ as $s\to 0$. In Eq. 73 we introduced the factor $s^{-1+\beta}$ rather than $s^{-2+\beta}$, which guarantees that $L_{0}(\alpha,\beta,\phi;s)$ vanishes at $s=0$, and subsequently yields the numerically small (in absolute value) lower bound $m_{0}$ as shown. The price is that in the error bounds of Lemma 3.7 below there is a factor $n^{-1+\alpha}$ rather than a smaller term $n^{-2+\alpha}$, but this weaker decay is more than sufficient for $n\geq 5$ which we consider.
Likewise, we use the factor $s^{-2-\alpha}$ as opposed $s^{-1-\alpha}$ in the definition Eq. 75 (cf. Eq. 71) for the same reason. On the set set $\alpha,\beta\in[0,1]$, $\phi\in[\phi_{0},\pi]$ and $s\in[1,\infty)$ the function $L_{\infty}(\alpha,\beta,\phi;s)$ is continuous, bounded and $\mathcal{O}(s^{-1})$ as $s\to\infty$.
See Figs. 1 and 2 for graphs of $L_{0}(\alpha,\beta,\phi_{0};s_{0})$ and $L_{\infty}(\alpha,\beta,\phi_{0};s_{\infty})$ for $\alpha,\beta\in[0,1]$. In [Dunster:2026:BWM] we provide plots of $L_{0}(\alpha,\beta,\phi_{0};s)$ and $L_{0}(\alpha,\beta,\pi;s)$ animated for $s\in[0,1]$, and $L_{\infty}(\alpha,\beta,\phi_{0};s)$ and $L_{\infty}(\alpha,\beta,\pi;s)$ animated for $s\in[1,30]$.
Figure 1: Graph of $L_{0}(\alpha,\beta,\phi_{0};s_{0})$ for $\alpha,\beta\in[0,1]$.
Figure 2: Graph of $L_{\infty}(\alpha,\beta,\phi_{0};s_{\infty})$ for $\alpha,\beta\in[0,1]$.
Before discussing the computations leading to Eqs. 76 and 77, we apply these values to obtain the following desired explicit lower bounds.
Lemma 3.7.
The error terms in Eq. 54 satisfy
(78) $\mathcal{E}_{0}(\alpha,\beta,\phi;n)>m_{0}\,\Gamma(2-\beta)\,n^{-1+\alpha},$
and
(79) $\mathcal{E}_{\infty}(\alpha,\beta,\phi;n)\geq m_{\infty}\,\Gamma(3+\alpha,n)\,n^{-2-\beta}.$
uniformly for $n>0$, $\alpha,\beta\in[0,1]$ and $\phi\in[\phi_{0},\pi]$, where $\phi_{0}=0.061$.
Proof 3.8.
By the definition Eq. 72 we have under the hypotheses of the lemma
$L(\alpha,\beta,\phi;s)\;\geq\;m_{0}\,s^{1-\beta},$
Inserting this into Eq. 57 and using the elementary inequality
(80) $\int_{a}^{b}f(s)\,g(s)\,ds\;\geq\;f_{0}\int_{a}^{b}g(s)\,ds\quad(f(s)\geq f_{0},\,g(s)\geq 0),$
yields, since $m_{0}<0$,
$\mathcal{E}_{0}(\alpha,\beta,\phi;n)\geq m_{0}\,n^{1+\alpha-\beta}\int_{0}^{1}s^{1-\beta}e^{-ns}\,ds>m_{0}\,n^{1+\alpha-\beta}\int_{0}^{\infty}s^{1-\beta}e^{-ns}\,ds,$
and Eq. 78 then follows from referring to Eq. 51.
Similarly, combining Eq. 58 with the pointwise lower bound
$L(\alpha,\beta,\phi;s)\;\geq\;m_{\infty}\,s^{2+\alpha},$
which comes from Eq. 74, and using the same inequality Eq. 80, we obtain for $\alpha,\beta\in[0,1]$ and $\phi\in[\phi_{0},\pi]$,
$\mathcal{E}_{\infty}(\alpha,\beta,\phi;n)\geq m_{\infty}\,n^{1+\alpha-\beta}\int_{1}^{\infty}s^{2+\alpha}e^{-ns}\,ds.$
With the change of variable $s\to s/n$ and referring to the definition of the incomplete gamma function [NIST:DLMF, Eq. 8.2.2] we arrive at Eq. 79.
3.1 Numerical computation of $m_{0}$ and $m_{\infty}$
We minimized $L_{0}(\alpha,\beta,\phi;s)$ by a direct numerical search in Maple (Digits =60=60) over $\alpha,\beta\in[0,1]$ and $\phi\in[\phi_{0},\pi]$, with the ss–dependence handled as follows. We sampled $(\alpha,\beta,\phi)$ on a fine uniform grid (using explicit limiting cases on the boundary faces), with the base grids in $\alpha$ and $\beta$ chosen to be denser near 0 and $1$, and $\phi$ sampled uniformly on $\phi_{0},\pi]$. Each cell of the $(\alpha,\beta,\phi)$ grid was refined once by inserting two equally spaced interior points in each coordinate direction.
For each sampled triple $(\alpha,\beta,\phi)$, we regarded $L_{0}(\alpha,\beta,\phi;s)$ as a function of the single variable $s\in[0,1]$, and evaluated it at the endpoint $s=1$ and at any interior critical points $s\in(0,1)$ satisfying $\partial L_{0}(\alpha,\beta,\phi;s)/\partial s=0$. The critical points were located using fsolve on sign-change brackets, based on an exact expression for the $s$-derivative. Since $L_{0}(\alpha,\beta,\phi;0)=0$, the endpoint $s=0$ did not need to be sampled.
The boundary cases $\alpha=0,1$, $\beta=0,1$, and $\phi=\pi$ are removable singularities; these were evaluated stably using the exact limiting formulas in Eqs. 39, 40, 41, 42, 43, 59, 60, 61, 62, 63, 64, 65, and 73.
We carried out a similar grid search for the normalised kernel $L_{\infty}(\alpha,\beta,\phi;s)$ defined in Eq. 75, again in Maple with 60-digit arithmetic. We sampled on the same $\alpha$, $\beta$ and $\phi$ 3-D grid as before, but this time for $s$ at 1 and 30, and at any critical points within $(1,30)$. At each quadruple point we evaluated $L_{\infty}(\alpha,\beta,\phi;s)$, storing the current minimum. Over this truncated box, with $s$ taking only the boundary and critical interior values, the minimum recorded value was given by Eq. 77.
That we were able to evaluate $m_{\infty}$ in Eq. 74 by restricting our search for the infimum of $L_{\infty}(\alpha,\beta,\phi;s)$ to $1\leq s\leq 30$ is due to the following.
Proposition 3.9.
(81) $\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi),\,s\in[1,30]\end{subarray}}\ L_{\infty}(\alpha,\beta,\phi;s)=-0.03332478558\cdots\\ <\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi),\,s\in[30,\infty)\end{subarray}}L_{\infty}(\alpha,\beta,\phi;s)=-0.008293818653\cdots,$
with the latter infimum attained at at $\alpha=0$, $\beta=1$, $\phi=0.061$, and $s=30$.
To examine the large-$s$ regime numerically in a stable way, we re-expressed $L_{\infty}(\alpha,\beta,\phi;s)$ in terms of exponentials with negative exponent; for example,
$\left(\frac{s}{e^{s}-1}\right)^{\beta}\left\{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}\right\}\\ =\left(\frac{s}{1-e^{-s}}\right)^{\beta}e^{(\alpha-\beta)s}\left\{(1+e^{-s})^{\alpha}-\left(1+e^{-2s}-2e^{-s}\cos(\phi)\right)^{1/2}\right\},$
which appears in Eq. 75 via Eqs. 24, 32, and 59. We then minimized the resulting large $s$ stable form of $L_{\infty}(\alpha,\beta,\phi;s)$ over $\alpha,\beta\in(0,1)$, $\phi\in[\phi_{0},\pi)$ and $s\in[30,S]$ using Maple’s global search (Optimization[Minimize]) (using high precision arithmetic), for various values of $S\in[100,10^{5}]$. In every run, the optimizer returned the same minimum at the lower boundary $s=30$, with the value given in Eq. 81. This provides a practical numerical certification that the infimum defining $m_{\infty}$ is achieved within the finite window $1\leq s\leq 30$, bearing in mind, from Eqs. 71 and 75, that $L_{\infty}(\alpha,\beta,\phi;s)=\mathcal{O}(s^{-1})$ as $s\to\infty$. This is illustrated by our optimization computation for $s=10^{5}$ giving the value
(82) $\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi)\end{subarray}}L_{\infty}(\alpha,\beta,\phi;10^{5})=L_{\infty}(0,1,\phi_{0};10^{5})=-0.2665816964\times 10^{-5}.$
Finally, we also used Optimization[Minimize] to independently verify the grid results that yielded Eqs. 76 and $7$ for $s\in[0,1]$ and $s\in[1,30]$, respectively. The Maple source code for the evaluation of $m_{0}$ and $m_{\infty}$ is provided in [Dunster:2026:BWM].
4 Proof of positivity for $n=5$
Let $w_{n}(\alpha,\beta,x)$ be the scaled hypergeometric function defined by Eq. 12, $A_{1,2}(\alpha,\phi)$ be given by Eqs. 52 and 53, and $m_{0}$ and $m_{\infty}$ be the infima given by Eqs. 76 and 77. Then from Eq. 28, the scaling Eq. 55 together with Euler’s reflection identity Eq. 8, and the decomposition Eqs. 54 and 56 combined with the bounds Eqs. 78 and 79, we deduce the following.
Theorem 4.1.
For $n=1,3,5,\dots$, $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi)$
(83) $w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\geq Q_{n}(\alpha,\beta,\phi)\,P_{n}(\alpha,\beta,\phi),$
where
(84) $P_{n}(\alpha,\beta,\phi)=\frac{1}{\alpha+\beta}\left[A_{1}(\alpha,\phi)\,n^{\alpha}\left\{1+\frac{(\alpha-\beta)(1-\beta)}{2n}\right\}\right.\\ \left.-\frac{A_{2}(\alpha,\beta,\phi)}{n^{\beta}}\left\{1+\frac{(\alpha-\beta)(1+\alpha)}{2n}\right\}\right]\\ -\frac{\left|m_{0}\right|\,\Gamma(2-\beta)}{n^{1-\alpha}}-\frac{\left|m_{\infty}\right|\,\Gamma(3+\alpha,n)}{n^{2+\beta}},$
and
(85) $Q_{n}(\alpha,\beta,\phi)=\frac{\pi\,(\alpha+\beta)\sin(\pi\,\alpha)n!(\pi-\phi)^{2}}{(1-\alpha)\Gamma(\beta)\Gamma(n-\alpha)n^{1+\alpha-\beta}}.$
Note that at the boundaries
(86) $P_{n}(0,1,\phi)=\frac{1}{\pi(\pi-\phi)^{2}}\left[\ln\left\{\csc\left(\tfrac{1}{2}\phi\right)\right\}-\frac{\csc\left(\tfrac{1}{2}\phi\right)-1}{2n}\left(1-\frac{1}{2n}\right)\right]\\ -\frac{|m_{0}|}{n}-\frac{\left|m_{\infty}\right|}{n\,e^{n}}\left(1+\frac{2}{n}+\frac{2}{n^{2}}\right),$
(87) $P_{n}(1,0,\phi)=n\left(2+\frac{1}{n}\right)\frac{1-\sin(\tfrac{1}{2}\phi)}{\pi\,(\pi-\phi)^{2}}-\left|m_{0}\right|-\frac{n\left|m_{\infty}\right|}{e^{n}}\left(1+\frac{3}{n}+\frac{6}{n^{2}}+\frac{6}{n^{3}}\right),$
(88) $P_{n}(1,1,\phi)=\frac{n\left\{1-\sin(\tfrac{1}{2}\phi)\right\}}{\pi\,(\pi-\phi)^{2}}-\left|m_{0}\right|-\frac{\left|m_{\infty}\right|}{e^{n}}\left(1+\frac{3}{n}+\frac{6}{n^{2}}+\frac{6}{n^{3}}\right),$
and
(89) $P_{n}(0,0,\phi)=\frac{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}{\pi(\pi-\phi)^{2}}\left[\ln(n)+\gamma-\frac{1}{2}\ln\left\{\frac{\csc(\tfrac{1}{2}\phi)}{4}\right\}\right]\\ -\frac{|m_{0}|}{n}-\frac{\left|m_{\infty}\right|}{e^{n}}\left(1+\frac{2}{n}+\frac{2}{n^{2}}\right).$
All the above can easily be shown to be positive for $n\geq 5$ and $\phi\in[\phi_{0},\pi)$, with finite positive limits as $\phi\to\pi$.
In addition, $P_{n}(\alpha,\beta,\pi)$ is given by Eq. 84 with $A_{1}(\alpha,\pi)$ and $A_{2}(\alpha,\beta,\pi)$ being given by Eqs. 62 and 65, and incidentally from this (for $n=5$) we find numerically that
(90) $\inf_{\alpha,\beta\in(0,1)}P_{5}(\alpha,\beta,\pi)=P_{5}(0,1,\pi)=\frac{9}{80\pi}-\frac{|m_{0}|}{5}-\frac{37\,|m_{\infty}|}{125\,e^{5}}=0.01621575275\cdots.$
For $n=5$ the desired positivity of the infimum is given by the following.
Proposition 4.2.
For $\phi_{0}=0.061$
(91) $\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi)\end{subarray}}P_{5}(\alpha,\beta,\phi)=P_{5}(0,1,\phi_{0})\\ =\frac{1}{\pi(\pi-\phi_{0})^{2}}\left[\ln\left\{\csc\left(\tfrac{1}{2}\phi_{0}\right)\right\}-\frac{9\left\{\csc\left(\tfrac{1}{2}\phi_{0}\right)-1\right\}}{100}\right]\\ -\frac{|m_{0}|}{5}-\frac{37\left|m_{\infty}\right|}{125\,e^{5}}=0.001500310752\cdots.$
Figure 3: Graph of $P_{5}(\alpha,\beta,\phi_{0})$ for $\alpha,\beta\in[0,1]$.
Figure 4: Graph of $P_{5}(\alpha,\beta,\phi_{0})$ for $\alpha\in[0,0.1]$ and $\beta\in[0.9,1]$.
Remark 4.3.
See Figs. 3 and 4 for surface plots of $P_{5}(\alpha,\beta,\phi_{0})$ (the latter depicting the surface close to the minimum). That the (positive) infimum Eq. 91 is relatively small is a consequence of our choice of $\phi_{0}=0.061$ being as small as possible. We chose the smallest such value (to three digits) that would still result in Eq. 91 being positive. We see from Eq. 53 that $A_{2}(\alpha,\beta,\phi)\to-\infty$ as $\phi\to 0^{+}$. In addition, the infima $m_{0}$ and $m_{\infty}$, as given by Eqs. 59, 72, 73, 74, and 75, also depend on $\phi_{0}$, and they too approach -$\infty$ as $\phi_{0}\to 0^{+}$. Hence from Eq. 84 $P_{5}(\alpha,\beta,\phi)\to$-$\infty$ in this limit (noting from Eq. 52 that $A_{1}(\alpha,\beta,\phi)$ remains bounded).
To numerically obtain the infimum of $P_{5}(\alpha,\beta,\phi)$ given in Eq. 84 we carried out a 2-D $\alpha,\beta$ grid search in Maple (using high-precision arithmetic), storing at each step the smallest current value of $P_{5}(\alpha,\beta,\phi)$. We fixed $\phi_{0}=0.061$ and $n=5$, and used the numerical values of $m_{0}$ and $m_{\infty}$ from Eqs. 76 and 77. The search sampled $\alpha$ and $\beta$ over $60^{2}$ point pairs on a grid biased toward 0 and $1$ (including boundary values).
For each grid pair $(\alpha,\beta)$ we minimized over $\phi\in[\phi_{0},\pi]$ by evaluating $P_{5}(\alpha,\beta,\phi)$ only at the endpoints $\phi=\phi_{0}$ and $\phi=\pi$, together with any interior stationary points in $(\phi_{0},\pi)$ where $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi=0$. These stationary points were computed from exact closed-form formulas for the derivatives, using fsolve with sign-change bracketing. Since the cases $\alpha,\beta\in\{0,1\}$ and $\phi=\pi$ are removable singularities, near-edge values were snapped to $\alpha,\beta=0$ or $1$, and $\phi=\pi$, with both $P_{5}(\alpha,\beta,\phi)$ and $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi$ being evaluated via the corresponding exact limiting formulas on edges and corners (see Eqs. 62, 65, 86, 87, 88, 89, and 90). Over the resulting candidate set in $\phi$, the smallest value recorded on the $(\alpha,\beta)$ grid yielded Eq. 91, attained as stated at $(\alpha,\beta,\phi)=(0,1,\phi_{0})$. The exact value shown comes from Eq. 87.
Again, as an independent check, we used the Maple routine Minimize over the box $\alpha,\beta\in[0,1]$, $\phi\in[\phi_{0},\pi]$, with Digits set very high, to verify Eq. 91. Finally, in [Dunster:2026:BWM] surface plots of $P_{5}(\alpha,\beta,\phi)$ for $\alpha,\beta\in[0,1]$, as well as for $(\alpha,\beta)$ near $(0,1)$, animated over $\phi_{0}\leq\phi\leq\pi$, provide further visual confirmation. Maple source code for the computation of Eq. 91 is also given.
5 Proof for $n=7,9,11,13,\ldots$
Define
(92) $\nu=\alpha+\beta\in(0,2),$
(93) $c_{1}(\alpha,\beta)=\tfrac{1}{2}(\alpha-\beta)(1-\beta),$
(94) $c_{2}(\alpha,\beta)=\tfrac{1}{2}(\alpha-\beta)(1+\alpha),$
and
(95) $J(\alpha,\beta,\phi;n)=n^{1-\alpha}P_{n}(\alpha,\beta,\phi),$
so that, using Eq. 84,
(96) $J(\alpha,\beta,\phi;n)=\nu^{-1}A_{1}(\alpha,\phi)\left\{n+c_{1}(\alpha,\beta)\right\}\\ -\nu^{-1}n^{-\nu}A_{2}(\alpha,\beta,\phi)\left\{n+c_{2}(\alpha,\beta)\right\}-|m_{0}|\,\Gamma(2-\beta)-|m_{\infty}|\,g_{\alpha,\beta}(n),$
where
(97) $g_{\alpha,\beta}(n)=\Gamma(3+\alpha,n)\,n^{-1-\nu}.$
The purpose of this section is to prove the following, which in conjunction with Eqs. 83, 91, and 95, proves that $w_{n}(\alpha,\beta,-1)-|w_{n}(\alpha,\beta,e^{i\phi})|\geq 0$ for $\phi\in[\phi_{0},\pi]$ and $n=7,9,11,\ldots$.
Theorem 5.1.
For $n\geq 5$ (not necessarily integral), $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi)$
(98) $J(\alpha,\beta,\phi;n)\geq J(\alpha,\beta,\phi;5)>0.$
From Proposition 4.2 and Eq. 95 it suffices to show that $\partial J/\partial n\geq 0$ for all $n\geq 5$, where $n$ is temporarily regarded as a continuous real variable rather than an integer. To this end, on differentiating Eq. 96 with respect to $n$, we find
(99) $\partial J(\alpha,\beta,\phi;n)/\partial n=\nu^{-1}A_{1}(\alpha,\phi)-\nu^{-1}(1-\nu)A_{2}(\alpha,\beta,\phi)\,n^{-\nu}\\ +A_{2}(\alpha,\beta,\phi)\,c_{2}(\alpha,\beta)\,n^{-1-\nu}-|m_{\infty}|\,g_{\alpha,\beta}^{\prime}(n).$
Thus $\partial J/\partial n\sim\nu^{-1}A_{1}>0$ for large $n$ and $\nu>0$, which indicates, at least asymptotically, the desired positivity, and hence that the desired bound Eq. 98 holds. We now proceed to confirm this.
To do so, we note from Eq. 94 that $c_{2}(\alpha,\beta)$, which appears in the third term on the RHS of Eq. 99, can take either sign, and so we split the analysis into the two cases $c_{2}(\alpha,\beta)\geq 0$ and $c_{2}(\alpha,\beta)<0$.
5.1 Case 1: $c_{2}(\alpha,\beta)\geq 0\;(\alpha\geq\beta)$
Firstly, from Eq. 97 and using $\partial\Gamma(a,x)/\partial x=-x^{-1+a}e^{-x}$ yields
(100) $g_{\alpha,\beta}^{\prime}(n)=-(1+\nu)\,n^{-2-\nu}\,\Gamma(3+\alpha,n)-n^{1-\beta}e^{-n}<0,$
for $n>0$ and $\alpha,\beta\in(0,1]$.
Now for the present case, the $c_{2}(\alpha,\beta)$ term in Eq. 99 is nonnegative, and by Eq. 100 the last term is also nonnegative. Consequently, noticing that $n^{-\nu}\leq 5^{-\nu}$ for $n\geq 5$, we have
(101) $\partial J(\alpha,\beta,\phi;n)/\partial n\geq F_{1}(\alpha,\beta,\phi)\quad(n\geq 5,\,0\leq\beta\leq\alpha\leq 1),$
where
(102) $F_{1}(\alpha,\beta,\phi)=\frac{1}{\nu}\left[A_{1}(\alpha,\phi)-\frac{\max\{1-\nu,0\}}{5^{\nu}}A_{2}(\alpha,\beta,\phi)\right].$
Since $A_{1}(\alpha,\phi)$ and $A_{2}(\alpha,\beta,\phi)$ are smooth in $(\alpha,\beta,\phi)$ and from Eqs. 60 and 63
$A_{1}(\alpha,\phi)-A_{2}(\alpha,\beta,\phi)(1-\nu)5^{-\nu}=\mathcal{O}(\nu)\quad(\nu\to 0),$
the quantity $F_{1}(\alpha,\beta,\phi)$ is continuous and bounded on the parameter regime under consideration. Incidentally, since $A_{1}(\alpha,\phi)>0$ it is certainly clear that $F_{1}(\alpha,\beta,\phi)>0$ for $\nu=\alpha+\beta\geq 1$.
5.2 Case 2: $c_{2}(\alpha,\beta)<0\;(\alpha<\beta)$
For $n\geq 5$ we have
$A_{2}(\alpha,\beta,\phi)(1-\nu)\,n^{-\nu}\geq-A_{2}(\alpha,\beta,\phi)|1-\nu|\,5^{-\nu},$
and
$\nu A_{2}(\alpha,\beta,\phi)c_{2}(\alpha,\beta)\,n^{-1-\nu}\geq-\nu A_{2}(\alpha,\beta,\phi)|c_{2}(\alpha,\beta)|\,5^{-1-\nu}.$
Therefore, in conjunction with Eqs. 99 and 100, it is evident that
(103) $\partial J(\alpha,\beta,\phi;n)/\partial n\geq F_{2}(\alpha,\beta,\phi)\quad(n\geq 5,\,0\leq\alpha<\beta\leq 1),$
where
(104) $F_{2}(\alpha,\beta,\phi)=\frac{1}{\nu}\left[A_{1}(\alpha,\phi)-\frac{1}{5^{\nu}}\left\{|1-\nu|+\frac{\nu|c_{2}(\alpha,\beta)|}{5}\right\}A_{2}(\alpha,\beta,\phi)\right].$
Thus, in summary, to establish $\partial J/\partial n\geq 0$ for $n\geq 5$ it is sufficient from Eqs. 101 and 103 to check that $F_{1}(\alpha,\beta,\phi)\geq 0$ for $0\leq\beta\leq\alpha\leq 1$, and $F_{2}(\alpha,\beta,\phi)\geq 0$ for $0\leq\alpha<\beta\leq 1$. The following then completes the proof of Theorem 5.1.
Proposition 5.2.
For $\phi_{0}=0.061$ we have
(105) $\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1),\,\alpha\geq\beta\\ \phi\in[\phi_{0},\pi)\end{subarray}}F_{1}(\alpha,\beta,\phi)=0.03251857515\cdots,$
attained at $(\alpha,\beta,\phi)=(1,1,\phi_{0})$, and
(106) $\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1),\,\alpha<\beta\\ \phi\in[\phi_{0},\pi)\end{subarray}}F_{2}(\alpha,\beta,\phi)=0.03204047407\cdots,$
attained at $(\alpha,\beta,\phi)=(\alpha_{2},1,\phi_{0})$, where $\alpha_{2}=0.8583872779\cdots$. In particular, both infima are strictly positive.
The computation of these values was performed in a similar manner to the method used for Proposition 4.2, that is, sampling over a fine $\alpha,\beta$ grid, and at $\phi$ boundary points ($\phi_{0}$ and $\pi$) and at any interior critical points. The infima were then independently confirmed using the Maple Optimization[Minimize] package. The grid and optimization code can be found in [Dunster:2026:BWM].
For the boundary points we again utilised exact limits, which in this case are given by
(107) $F_{1}(0,0,\phi)=F_{2}(0,0,\phi)=\frac{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}\left[2\{1+\gamma+\ln(10)\}+\ln\left\{\sin(\tfrac{1}{2}\phi)\right\}\right]}{2\pi\,\Delta^{2}},$
(108) $F_{1}(1,1,\phi)=\frac{1-\sin\left(\tfrac{1}{2}\phi\right)}{\pi(\pi-\phi)^{2}},$
and
(109) $F_{2}(\alpha,1,\phi)=\frac{1}{1+\alpha}\left[A_{1}(\alpha,\phi)-\frac{(1-\alpha)\,C_{2}(1,\phi)}{5^{1+\alpha}\,\pi}\left\{\alpha+\frac{1}{10}(1+\alpha)^{2}(1-\alpha)\right\}\right],$
where from Eqs. 23 and 31
(110) $C_{2}(1,\phi)=\frac{\csc\left(\tfrac{1}{2}\phi\right)-1}{2(\pi-\phi)^{2}}.$
In addition
(111) $F_{1}(\alpha,\beta,\pi)=\frac{1}{\nu}\left[A_{1}(\alpha,\pi)-\frac{\max\{1-\nu,0\}A_{2}(\alpha,\beta,\pi)}{5^{\nu}}\right],$
(112) $F_{2}(\alpha,\beta,\pi)=\frac{1}{\nu}\left[A_{1}(\alpha,\pi)-\frac{A_{2}(\alpha,\beta,\pi)}{5^{\nu}}\left(|1-\nu|+\frac{\nu\left|c_{2}(\alpha,\beta)\right|}{5}\right)\right],$
where $A_{1}(\alpha,\pi)$, $A_{2}(\alpha,\beta,\pi)$ and $c_{2}(\alpha,\beta)$ are given by Eqs. 62, 65, and 94.
Incidentally, in comparison to Eqs. 105 and 106, from Eqs. 111 and 112 we find
$\inf_{\alpha\geq\beta}F_{1}(\alpha,\beta,\pi)=F_{1}(1,1,\pi)=\frac{1}{8\pi}>\inf_{\alpha<\beta}F_{2}(\alpha,\beta,\pi)=0.03287693288\cdots,$
the latter being attained at $\alpha=0.4852393602\cdots$ and $\beta=1$.
Figure 5: Graph of $F(\alpha,\beta,\phi_{0})$ for $\alpha,\beta\in[0,1]$.
In Fig. 5 the surface of $F(\alpha,\beta,\phi_{0})$ is shown, where
(113) $F(\alpha,\beta,\phi)=\begin{cases}F_{1}(\alpha,\beta,\phi)&(\alpha\geq\beta)\\[4.30554pt] F_{2}(\alpha,\beta,\phi)&(\alpha<\beta)\end{cases}\;,$
and this illustrates the infima of Proposition 5.2. For an animated version over $\phi\in[\phi_{0},\pi]$ see [Dunster:2026:BWM].
6 Extension to $\phi\in[0,\phi_{0})$ ($\theta\in(\pi-\phi_{0},\pi]$)
The Watson large $n$ expansion breaks down near $\phi=0$, as can be seen from Eq. 31 which becomes unbounded in this limit. A possible approach for $\phi\in[0,\phi_{0})$ ($\phi_{0}=0.061$) is to replace Eq. 35 with a uniform approximation for large $n$, valid near $\phi=0$, involving the Kummer function [NIST:DLMF, Eq. 13.4.4]
(114) $U(p,q,z)=\frac{1}{\Gamma(p)}\int_{0}^{\infty}e^{-zt}\,t^{p-1}(1+t)^{q-p-1}dt\quad(\Re(p)>0,\,|\arg(z)|<\tfrac{1}{2}\pi).$
To see this, consider the troublesome second integral in Eq. 15, namely
(115) $\int_{0}^{\infty}(e^{s}-1)^{\alpha}(e^{s}-e^{-i\phi})^{-\beta}e^{-ns}\,ds,$
where $x=e^{i\phi}$. Again the factor $e^{-ns}$ localises the main contribution to small $s$ when $n$ is large. Thus, we use the first-order approximation of the integrand, namely
(116) $(e^{s}-1)^{\alpha}(e^{s}-e^{-i\phi})^{-\beta}\sim s^{\alpha}\,(s+a)^{-\beta}\quad(s\to 0),$
where
(117) $a=1-e^{-i\phi}=2i\,e^{-i\phi/2}\sin\left(\tfrac{1}{2}\phi\right)\sim i\phi\quad(\phi\to 0).$
Substituting Eq. 116 into Eq. 115 yields the formal approximation for large $n$
(118) $\int_{0}^{\infty}(e^{s}-1)^{\alpha}(e^{s}-e^{-i\phi})^{-\beta}e^{-ns}\,ds\sim\int_{0}^{\infty}s^{\alpha}(s+a)^{-\beta}e^{-ns}ds,$
and on comparing with Eq. 114 the right-hand side is explicitly given by
(119) $a^{1+\alpha-\beta}\Gamma(1+\alpha)U(1+\alpha,2+\alpha-\beta,na).$
In particular, since $a\sim i\phi$ as $\phi\to 0$, the Kummer argument $na$ captures the coupled small $\phi$ and large $n$ transitions for this second integral, which the simpler Watson expansion does not. For more details on this technique, see [Temme:2015:AMF, Chap. 28].
We finish by noting the following simple result, which confirms without using Kummer functions that Eq. 5 holds at $\theta=\pi$ ($\phi=0$) provided $\beta\geq\alpha$.
Lemma 6.1.
Let $0<\alpha\leq\beta<1$ and let $1\leq n\in\mathbb{N}$ be odd. Then
(120) $A_{n}(\alpha,\beta,1)>\bigl|A_{n}(\alpha,\beta,-1)\bigr|.$
Proof 6.2.
For $\omega=-1$ it is straightforward to show from Eq. 1 that
(121) $A_{n}(\alpha,\beta,-1)=\frac{\Gamma(n-\alpha+\beta)}{\Gamma(\beta-\alpha)\,n!}\geq 0\quad(0<\alpha\leq\beta<1),$
so $\bigl|A_{n}(\alpha,\beta,-1)\bigr|=A_{n}(\alpha,\beta,-1)$ under the hypothesis of the lemma.
Next, again from the generating function Eq. 1, expanding $(1+\omega z)^{\alpha}$ and $(1-z)^{-\beta}$ and collecting coefficients gives
(122) $A_{n}(\alpha,\beta,\omega)=\sum_{k=0}^{n}\binom{\alpha}{k}\,\frac{\Gamma(n-k+\beta)}{\Gamma(\beta)\,\Gamma(n-k+1)}\,\omega^{k},$
where $\binom{\alpha}{0}=1$ and, from Eq. 8,
(123) $\binom{\alpha}{k}=\frac{\Gamma(1+\alpha)}{\Gamma(1-k+\alpha)\,\Gamma(k+1)}\\ =(-1)^{k+1}\frac{\Gamma(1+\alpha)\,\Gamma(k-\alpha)\sin(\pi\alpha)}{\pi k!}\quad(k=1,2,3,\ldots n).$
Thus for $\alpha\in(0,1)$ the $k=0$ and the $k$-odd coefficients in Eq. 122 are positive. Therefore, using Eq. 122 with $\omega=\pm 1$ and the fact that $n$ is assumed to be odd, we obtain for $0<\alpha\leq\beta<1$
$A_{n}(\alpha,\beta,1)-A_{n}(\alpha,\beta,-1)=A_{n}(\alpha,\beta,1)-\left|A_{n}(\alpha,\beta,-1)\right|\\ =\sum_{k=0}^{n}\binom{\alpha}{k}\,\frac{\Gamma(n-k+\beta)}{\Gamma(\beta)\,\Gamma(n-k+1)}\left\{1-(-1)^{k}\right\}=2\sum_{\begin{subarray}{c}1\leq k\leq n\\ k\ \mathrm{odd}\end{subarray}}\!\!\binom{\alpha}{k}\,\frac{\Gamma(n-k+\beta)}{\Gamma(\beta)\,\Gamma(n-k+1)}>0.$
Acknowledgement
Financial support from Ministerio de Ciencia e Innovación project PID2024-159583NB-I00 (MICIU/ AEI / 10.13039/501100011033 / FEDER, UE) is acknowledged.
Conflict of interest
The author declares no conflicts of interest.

### v2

@@
 Eq. 2 and [NIST:DLMF, Eqs. 5.2.5, 5.2.6, 15.8.6] we obtain for positive integer $n$
-(9) $A_{n}(\alpha,\beta,\omega)=(-1)^{n+1}\frac{\Gamma(1+\alpha)\Gamma(n-\alpha)\sin(\pi\alpha)}{\pi n!}\\ \times\omega^{n}\ {}_{2}F_{1}\left(\beta,-n;\alpha-n+1;-1/\omega\right).$
+(9) $A_{n}(\alpha,\beta,\omega)=\frac{(-1)^{n+1}}{\pi}\omega^{n}w_{n}(\alpha,\beta,-1/\omega),$
+where
+(10) $w_{n}(\alpha,\beta,x)=\frac{1}{n!}\Gamma(n-\alpha)\Gamma(1+\alpha)\sin(\pi\alpha)\,{}_{2}F_{1}\!\left(\beta,-n;\alpha-n+1;x\right).$
 Letting $x=-1/\omega=\exp(i\phi)$, so that
@@
 (12) $w_{n}(\alpha,\beta,-1)\geq\left|w_{n}(\alpha,\beta,e^{i\phi})\right|,$
-for $\alpha,\beta\in(0,1]$, $\phi\in[0,\pi]$ and $n=3,5,7,\ldots$, where
-(12) $w_{n}(\alpha,\beta,x)=\sin(\pi\alpha)\,{}_{2}F_{1}\!\left(\beta,-n;\alpha-n+1;x\right).$
-Note from Eqs. 1, 4, 9, and 12 that for $n=1,2,3,\ldots$
-(13) $\lim_{\alpha\to 0}w_{n}(\alpha,\beta,x)=-\frac{\pi\Gamma(\beta+n)}{\Gamma(\beta)(n-1)!}x^{n},$
+for $\alpha,\beta\in(0,1]$, $\phi\in[0,\pi]$ and $n=3,5,7,\ldots$.
+Note from Eqs. 1, 4, 9, and 10 that for $n=1,2,3,\ldots$
+(13) $\lim_{\alpha\to 0}w_{n}(\alpha,\beta,x)=-\frac{\pi\,\Gamma(\beta+n)}{\Gamma(\beta)\,n!}\,x^{n}.$
 and for $n=2,3,4\ldots$
-(14) $\lim_{\alpha\to 1}w_{n}(\alpha,\beta,x)=\frac{\pi\,\Gamma(\beta+n-1)}{\Gamma(\beta)(n-2)!}\left\{n-(\beta+n-1)x\right\}\,x^{n-1}.$
+(14) $\lim_{\alpha\to 1}w_{n}(\alpha,\beta,x)=\frac{\pi\,\Gamma(\beta+n-1)}{\Gamma(\beta)\,n!}\left\{n-(\beta+n-1)x\right\}\,x^{n-1}.$
 We wish to attack the problem using Watson’s lemma for large $n$.
@@
 Assume $\alpha,\beta\in[0,1]$, $1\leq n\in\mathbb{N}$, and $x\in\mathbb{C}\setminus[0,\infty)$. Then
-(15) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)}{n!}w_{n}(\alpha,\beta,x)\\ =(-1)^{n+1}\sin(\pi\beta)(-x)^{\,n-\alpha}\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,(e^{s}-x)^{\alpha}e^{-ns}\,ds\\ +\sin(\pi\alpha)(-x)^{-\beta}\int_{0}^{\infty}(e^{s}-1)^{\alpha}\left(e^{s}-1/x\right)^{-\beta}e^{-ns}ds.$
+(15) $w_{n}(\alpha,\beta,x)=(-1)^{n+1}\sin(\pi\beta)(-x)^{\,n-\alpha}\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,(e^{s}-x)^{\alpha}e^{-ns}\,ds\\ +\sin(\pi\alpha)(-x)^{-\beta}\int_{0}^{\infty}(e^{s}-1)^{\alpha}\left(e^{s}-1/x\right)^{-\beta}e^{-ns}ds.$
 When $\beta=1$ it is understood that the limit of the first term on the RHS applies.
 Proof 2.2.
-From [NIST:DLMF, Eqs. 15.1.2 and 15.6.4]
-(16) $_{2}F_{1}(a,b;c;z)=\frac{e^{-b\pi i}\Gamma(1-b)\Gamma(c)}{2\pi i\,\Gamma(c-b)}\int_{1}^{(0+)}\frac{t^{\,b-1}(1-t)^{c-b-1}}{(1-zt)^{a}}\,dt,$
-with $b\neq 1,2,3,\dots$, $c\neq 0,-1,-2,\dots$, $\Re(c-b)>0$, and $|\arg(1-z)|<\pi$.
-Here the contour is the standard loop starting at $t=1$ winding once counterclockwise around $t=0$ and returning to $t=1$.
-So with $a=\beta$, b=nb=-n, $c=1+\alpha-n$ and $x=1/\xi$ we have from Eqs. 8 and 12 the integral representation
-(17) $w_{n}(\alpha,\beta,1/\xi)=\frac{i\,n!}{2\Gamma(n-\alpha)\Gamma(1+\alpha)}(-\xi)^{\beta}I(\xi),$
+From Eq. 1 and Cauchy’s integral formula, for any $c\in(0,1)$ we have for $|\omega|<1$
+(16) $A_{n}(\alpha,\beta,\omega)=\frac{1}{2\pi i}\oint_{|z|=c}\frac{(1+\omega z)^{\alpha}}{(1-z)^{\beta}}\,\frac{dz}{z^{n+1}},$
+with the contour positively orientated. Setting $\omega=-\xi$ and using Eq. 9 (with $x=1/\xi=-1/\omega$) yields
+(17) $w_{n}(\alpha,\beta,1/\xi)=-\frac{1}{2i\xi^{n}}\oint_{|z|=c}\frac{(1-\xi z)^{\alpha}}{(1-z)^{\beta}}\,\frac{dz}{z^{n+1}}.$
+Now we temporarily assume $\xi=e^{i\varphi}$ with $\varphi\in(0,\pi]$ and $\alpha,\beta\in(0,1)$, and make the change of variable $z=t/\xi$. As a result we obtain
+(18) $w_{n}(\alpha,\beta,1/\xi)=\tfrac{1}{2}i(-\xi)^{\beta}I(\xi),$
 where
-(18) $I(\xi)=\int_{1}^{(0+)}\frac{t^{-n-1}(1-t)^{\alpha}}{(t-\xi)^{\beta}}\,dt.$
-We temporarily assume $\xi=e^{i\theta}$ with $\theta\in(0,\pi]$ and $\alpha,\beta\in(0,1)$, and for these values the integral converges, $\xi$ lies off the cut $1,\infty)$, and the branch points $t=1$ and $t=\xi$ are distinct.
+(19) $I(\xi)=\oint_{|t|=c}\frac{t^{-n-1}(1-t)^{\alpha}}{(t-\xi)^{\beta}}\,dt,$
+with the contour again positively orientated and $c\in(0,1)$.
+Next, $\xi=e^{i\varphi}$ lies off the cut $1,\infty)$, and the branch points $t=1$ and $t=\xi$ are distinct.
 The contour initially lies inside the unit circle $|t|=1$ and so avoids the branch point $t=\xi$.
-We take the branch cut for $(1-t)^{\alpha}$ to be $1,\infty)$ (principal branch on $\mathbb{C}\setminus[1,\infty)$), and the branch cut for $(t-\xi)^{-\beta}$ to be the ray extending from $\xi$ to $\infty$ with $\arg(t)=\arg(\xi)=\theta$.
-Since $\alpha,\beta\in(0,1)$ the branch points at $t=1$ and $t=\xi$ are locally integrable, and as $t\to\infty$ the integrand is $\mathcal{O}(t^{-n-1+\alpha-\beta})$ so contributions from circular arcs vanish at infinity.
-Hence the loop in Eq.
-18 may be deformed to four rays, two on either side of both of the cuts.
-As a result we arrive at
+We take the branch cut for $(1-t)^{\alpha}$ to be $1,\infty)$ (principal branch on $\mathbb{C}\setminus[1,\infty)$), and the branch cut for $(t-\xi)^{-\beta}$ to be the ray extending from $\xi$ to $\infty$ with $\arg(t)=\arg(\xi)=\varphi$.
+Since we assume that $\alpha,\beta\in(0,1)$, the branch points at $t=1$ and $t=\xi$ are locally integrable, and as $t\to\infty$ the integrand is $\mathcal{O}(t^{-n-1+\alpha-\beta})$, so contributions from circular arcs vanish at infinity.
+Hence, the loop in Eq.
+19 may be deformed to four rays, two on either side of both cuts.
+Consequently, we arrive at
 (20) $I(\xi)=I_{\xi}^{+}(\xi)-I_{\xi}^{-}(\xi)+I_{1}^{+}(\xi)-I_{1}^{-}(\xi),$
-where $I_{\xi}^{\pm}(\xi)$ denotes the integral of $t^{-n-1}(1-t)^{\alpha}(t-\xi)^{-\beta}$ from $t=\xi$ to $t=\infty\exp(i\theta)$ with $\arg(\xi-t)=\arg(\xi)\mp\pi$, and $I_{1}^{\pm}(\xi)$ denotes the integral of the same integrand from $t=1$ to $t=\infty$ with $\arg(1-t)=\mp\pi$.
+where $I_{\xi}^{\pm}(\xi)$ denotes the integral of $t^{-n-1}(1-t)^{\alpha}(t-\xi)^{-\beta}$ from $t=\xi$ to $t=\infty\exp(i\varphi)$ with $\arg(\xi-t)=\arg(\xi)\mp\pi$, and $I_{1}^{\pm}(\xi)$ denotes the integral of the same integrand from $t=1$ to $t=\infty$ with $\arg(1-t)=\mp\pi$.
 For the first pair make the change of variable $t=\xi v$ with $1\leq v<\infty$, so that for $I_{\xi}^{\pm}(\xi)$
 $(t-\xi)^{-\beta}=(-\xi)^{-\beta}(1-v)^{-\beta}=e^{\pm i\pi\beta}(-\xi)^{-\beta}(v-1)^{-\beta}\quad(1\leq v<\infty).$
-Therefore, since $(1-t)^{\alpha}=(1-\xi v)^{\alpha}$ takes the same value for both integrals for $\theta\in(0,\pi)$, we obtain
+Therefore, since $(1-t)^{\alpha}=(1-\xi v)^{\alpha}$ takes the same value for both integrals for $\varphi\in(0,\pi)$, we obtain
 (21) $I_{\xi}^{+}(\xi)-I_{\xi}^{-}(\xi)=2i(-1)^{n}\sin(\pi\beta)(-\xi)^{-n-\beta}\int_{1}^{\infty}v^{-n-1}(v-1)^{-\beta}(1-\xi v)^{\alpha}\,dv.$
@@
 (22) $I_{1}^{+}(\xi)-I_{1}^{-}(\xi)=-2i\sin(\pi\alpha)\int_{1}^{\infty}v^{-n-1}(v-1)^{\alpha}(v-\xi)^{-\beta}dv.$
-Inserting Eqs. 20 and 21 into Eq. 19 yields
+Inserting Eqs. 21 and 22 into Eq. 20 yields
 (23) $I(\xi)=2i(-1)^{n}\sin(\pi\beta)(-\xi)^{-n-\beta}\int_{1}^{\infty}v^{-n-1}(v-1)^{-\beta}(1-\xi v)^{\alpha}\,dv\\ -2i\sin(\pi\alpha)\int_{1}^{\infty}v^{-n-1}(v-1)^{\alpha}(v-\xi)^{-\beta}dv.$
@@
 Then, on using $(1-\xi e^{s})^{\alpha}=(-\xi)^{\alpha}(e^{s}-\xi^{-1})^{\alpha}$ and replacing $\xi$ by $1/x$, we confirm that Eq.
-15 holds for $\alpha,\beta\in(0,1)$ and $\arg(1/x)=\theta\in(0,\pi]$.
-Finally, the identity extends to $\alpha,\beta\in[0,1]$ by continuity (all boundary singularities are removable, with the case $\beta=1$ interpreted by the limiting value of the first term), and it extends from $x=e^{-i\theta}$, $\theta\in(0,\pi]$, to all $x\in\mathbb{C}\setminus[0,\infty)$ by analytic continuation in $x$.
+15 holds for $\alpha,\beta\in(0,1)$ and $\arg(1/x)=\varphi\in(0,\pi]$.
+Finally, the identity extends to $\alpha,\beta\in[0,1]$ by continuity (all boundary singularities are removable, with the case $\beta=1$ interpreted by the limiting value of the first term), and it extends from $x=e^{-i\varphi}$, $\varphi\in(0,\pi]$, to all $x\in\mathbb{C}\setminus[0,\infty)$ by analytic continuation in $x$.
 Our next step is to use this representation to give a suitable lower bound for the difference of the two terms in Eq. 12.
@@
 (26) $I_{1}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{-\beta}\left\{\left(e^{s}+1\right)^{\alpha}-R(s,\phi)^{\alpha}\right\}e^{-ns}ds,$
-(26) $I_{2}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{\alpha}\left\{\left(1+e^{s}\right)^{-\beta}-R(s,\phi)^{-\beta}\right\}e^{-ns}ds,$
+(27) $I_{2}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{\alpha}\left\{\left(e^{s}+1\right)^{-\beta}-R(s,\phi)^{-\beta}\right\}e^{-ns}ds,$
 and
@@
 Then
-(28) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}}{n!\Delta^{2}}\,\geq h(\alpha,\beta,\phi;n).$
+(29) $\Delta^{-2}\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}\geq h(\alpha,\beta,\phi;n).$
 Remark 2.4.
@@
 In Eq. 15 set x=1x=-1 and $x=\exp(i\phi)$, taking absolute values and then the integral triangle inequality on the latter. On subtraction of the two we deduce that
-(29) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)}{n!}\,\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}\\ \geq\sin(\pi\beta)\,\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,\left\{(e^{s}+1)^{\alpha}-\left|e^{s}-e^{i\phi}\right|^{\alpha}\right\}e^{-ns}\,ds\\ \quad+\sin(\pi\alpha)\,\int_{0}^{\infty}(e^{s}-1)^{\alpha}\,\left\{(e^{s}+1)^{-\beta}-\left|e^{s}-e^{-i\phi}\right|^{-\beta}\right\}e^{-ns}\,ds,$
+(30) $w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\\ \geq\sin(\pi\beta)\,\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,\left\{(e^{s}+1)^{\alpha}-\left|e^{s}-e^{i\phi}\right|^{\alpha}\right\}e^{-ns}\,ds\\ \quad+\sin(\pi\alpha)\,\int_{0}^{\infty}(e^{s}-1)^{\alpha}\,\left\{(e^{s}+1)^{-\beta}-\left|e^{s}-e^{-i\phi}\right|^{-\beta}\right\}e^{-ns}\,ds,$
 and Eq. 29 then follows on division of both sides by $\Delta^{2}$.
@@
 Let
-(30) $C_{1}(\alpha,\phi)=\frac{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}{\Delta^{2}},$
-(31) $C_{2}(\beta,\phi)=\frac{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}{\Delta^{2}},$
-(32) $K_{1}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{s}{e^{s}-1}\right)^{\beta}\frac{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}}{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}-1-\frac{\alpha-\beta}{2}\,s\right],$
+(31) $C_{1}(\alpha,\phi)=\Delta^{-2}\left[2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}\right],$
+(32) $C_{2}(\beta,\phi)=\Delta^{-2}\left[\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}\right],$
+(33) $K_{1}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{s}{e^{s}-1}\right)^{\beta}\frac{(e^{s}+1)^{\alpha}-R(s,\phi)^{\alpha}}{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}-1-\frac{\alpha-\beta}{2}\,s\right],$
 and
-(33) $K_{2}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{e^{s}-1}{s}\right)^{\alpha}\frac{R(s,\phi)^{-\beta}-(1+e^{s})^{-\beta}}{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}-1-\frac{\alpha-\beta}{2}\,s\right],$
+(34) $K_{2}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{e^{s}-1}{s}\right)^{\alpha}\frac{R(s,\phi)^{-\beta}-(e^{s}+1)^{-\beta}}{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}-1-\frac{\alpha-\beta}{2}\,s\right],$
 with $R(s,\phi)$ defined by Eq. 25. Then, for $\alpha,\beta\in(0,1)$ and $\phi\in[\phi_{0},\pi)$ for fixed $\phi_{0}>0$,
@@
 and
-(41) $K_{1}(0,0,\phi;s)=K_{2}(0,0,\phi;s)=\frac{1}{s^{2}}\left[\frac{\ln\left\{\left(1+e^{s}\right)/R(s,\phi)\right\}}{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}-1\right]\quad(0<\phi<\pi).$
+(42) $K_{1}(0,0,\phi;s)=K_{2}(0,0,\phi;s)=\frac{1}{s^{2}}\left[\frac{\ln\left\{\left(e^{s}+1\right)/R(s,\phi)\right\}}{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}-1\right]\quad(0<\phi<\pi).$
 Furthermore,
@@
 The coefficients $C_{1}(\alpha,\phi)$ and $C_{2}(\beta,\phi)$ defined in Eqs.
-30 and 31 remain smooth and strictly positive for $\phi\in[\phi_{0},\pi)$ and $\alpha,\beta\in(0,1)$, and have finite nonzero limits as $\phi\to\pi$; in particular
+31 and 32 remain smooth and strictly positive for $\phi\in[\phi_{0},\pi)$ and $\alpha,\beta\in(0,1)$, and have finite nonzero limits as $\phi\to\pi$; in particular
 (45) $C_{1}(\alpha,\phi)=\frac{2^{\alpha}\,\alpha}{8}+\mathcal{O}\left\{(\phi-\pi)^{2}\right\},$
@@
 Proof 3.3.
-From Eqs. 25 and 26
+From Eqs. 26 and 27
 (47) $I_{j}(\alpha,\beta,\phi;n)=\int_{0}^{\infty}G_{j}(\alpha,\beta,\phi;s)e^{-ns}ds\quad(j=1,2),$
 where
-(47) $G_{1}(\alpha,\beta,\phi;s)=\frac{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}}{\Delta^{2}(e^{s}-1)^{\beta}},$
+(48) $G_{1}(\alpha,\beta,\phi;s)=\frac{(e^{s}+1)^{\alpha}-R(s,\phi)^{\alpha}}{\Delta^{2}(e^{s}-1)^{\beta}},$
 and
-(48) $G_{2}(\alpha,\beta,\phi;s)=\frac{(e^{s}-1)^{\alpha}}{\Delta^{2}}\left[(1+e^{s})^{-\beta}-R(s,\phi)^{-\beta}\right].$
-Using a Taylor expansion at $s=0$ and the definition Eqs. 30 and 31 one finds that, as $s\to 0$,
+(49) $G_{2}(\alpha,\beta,\phi;s)=\frac{(e^{s}-1)^{\alpha}}{\Delta^{2}}\left[(e^{s}+1)^{-\beta}-R(s,\phi)^{-\beta}\right].$
+Using a Taylor expansion at $s=0$ and the definition Eqs. 31 and 32 one finds that, as $s\to 0$,
 $s^{\beta}G_{1},\,s^{-\alpha}G_{2}=\pm C_{1,2}\left\{1+\tfrac{1}{2}(\alpha-\beta)s+\mathcal{O}\left(s^{2}\right)\right\},$
@@
 (51) $G_{2}(\alpha,\beta,\phi;s)=-C_{2}(\beta,\phi)s^{\alpha}\left\{1+\tfrac{1}{2}(\alpha-\beta)s+s^{2}K_{2}(\alpha,\beta,\phi;s)\right\},$
-where the remainder kernels $K_{1,2}(\alpha,\beta,\phi;s)$ are given by Eqs. 32 and 33 and satisfy Eq. 38. Substituting Eqs. 49 and 50 into Eqs. 25 and 26, and using
+where the remainder kernels $K_{1,2}(\alpha,\beta,\phi;s)$ are given by Eqs. 33 and 34 and satisfy Eq. 39. Substituting Eqs. 50 and 51 into Eqs. 26 and 27, and using
 (52) $\int_{0}^{\infty}s^{\mu}e^{-ns}\,ds=\frac{\Gamma(1+\mu)}{n^{1+\mu}}\quad(\Re(n)>0,\Re(\mu)>-1),$
-we obtain Eqs. 34 and 35.
+we obtain Eqs. 35 and 36.
 Let us now consider the RHS of Eq. 29, suitably scaled. We introduce the prefactors
@@
 (54) $A_{2}(\alpha,\beta,\phi)=\frac{(1-\alpha)C_{2}(\beta,\phi)}{\sin(\pi\beta)\Gamma(1-\beta)}=\frac{(1-\alpha)\Gamma(\beta)\left[\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}\right]}{\pi(\pi-\phi)^{2}}.$
-Then, on combining Eqs. 27, 34, 35, 36, 37, 52, and 53, one arrives at:
+Then, on combining Eqs. 28, 35, 36, 37, 38, 53, and 54, one arrives at:
 Theorem 3.4.
@@
 60 $K_{1}(\alpha,\beta,\phi;s)$ and $K_{2}(\alpha,\beta,\phi;s)$ are given by Eqs.
-24, 32, and $3$.
-Note, from Eqs. 52 and 53,
+25, 33, and 34.
+Note, from Eqs. 53 and 54,
 (61) $A_{1}(0,\phi)=\frac{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}{\pi(\pi-\phi)^{2}},$
@@
 (66) $A_{2}(\alpha,\beta,\pi)=\frac{(1-\alpha)\Gamma(1+\beta)}{8\pi\,2^{\beta}},$
-and $L(\alpha,\beta,\pi;s)$ follows from Eqs. 39, 40, 59, 62, and 65.
+and $L(\alpha,\beta,\pi;s)$ follows from Eqs. 40, 41, 60, 63, and $6$.
 Furthermore, using the $\mathcal{O}(1)$ behaviour of $K_{1}(\alpha,\beta,\phi;s)$ and $K_{2}(\alpha,\beta,\phi;s)$ as $s\to 0$, one finds that
@@
 (69) $L(0,0,\pi;s)=\frac{1}{8\pi}\left[\mathrm{sech}^{2}(\tfrac{1}{2}s)\ln\left\{\coth(\tfrac{1}{2}s)\right\}+\ln(\tfrac{1}{2}s)\right]\\ =\frac{s^{2}\ln(s)}{32\pi}+\mathcal{O}\left(s^{2}\right)\quad(s\to 0).$
-We also note from Eqs. 23, 24, 32, 33, 59, 61, and 64 that
-(69) $L(0,1,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\frac{\csc(\tfrac{1}{2}\phi)-1}{2}\left(1-\frac{s}{2}\right)-\frac{1}{R(s,\phi)}+\frac{1}{1+e^{s}}\right\},$
+We also note from Eqs. 24, 25, 33, 34, 60, 62, and 65 that
+(70) $L(0,1,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\frac{\csc(\tfrac{1}{2}\phi)-1}{2}\left(1-\frac{s}{2}\right)-\frac{1}{R(s,\phi)}+\frac{1}{e^{s}+1}\right\},$
 (71) $L(1,0,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\sin(\tfrac{1}{2}\phi)(s+2)-R(s,\phi)+e^{s}-s-1\right\},$
-and from Eqs. 42, 43, and 59
+and from Eqs. 43, 44, and 60
 (72) $L(\alpha,\beta,\phi;s)\sim\frac{\alpha-\beta}{2(\alpha+\beta)}\left\{\frac{A_{2}(\alpha,\beta,\phi)}{\Gamma(1+\alpha)}\,s^{1+\alpha}-\frac{A_{1}(\alpha,\phi)}{\Gamma(1-\beta)}\,s^{1-\beta}\right\}\quad(s\to\infty).$
@@
 Eq.
-71) for the same reason.
+72) for the same reason.
 On the set set $\alpha,\beta\in[0,1]$, $\phi\in[\phi_{0},\pi]$ and $s\in[1,\infty)$ the function $L_{\infty}(\alpha,\beta,\phi;s)$ is continuous, bounded and $\mathcal{O}(s^{-1})$ as $s\to\infty$.
@@
 Figure 2: Graph of $L_{\infty}(\alpha,\beta,\phi_{0};s_{\infty})$ for $\alpha,\beta\in[0,1]$.
-Before discussing the computations leading to Eqs. 76 and 77, we apply these values to obtain the following desired explicit lower bounds.
+Before discussing the computations leading to Eqs. 77 and 78, we apply these values to obtain the following desired explicit lower bounds.
 Lemma 3.7.
@@
 The boundary cases $\alpha=0,1$, $\beta=0,1$, and $\phi=\pi$ are removable singularities; these were evaluated stably using the exact limiting formulas in Eqs.
-39, 40, 41, 42, 43, 59, 60, 61, 62, 63, 64, 65, and 73.
+40, 41, 42, 43, 44, 60, 61, 62, 63, 64, 65, 66, and 74.
 We carried out a similar grid search for the normalised kernel $L_{\infty}(\alpha,\beta,\phi;s)$ defined in Eq.
-75, again in Maple with 60-digit arithmetic.
+76, again in Maple with 60-digit arithmetic.
 We sampled on the same $\alpha$, $\beta$ and $\phi$ 3-D grid as before, but this time for $s$ at 1 and 30, and at any critical points within $(1,30)$.
@@
 Over this truncated box, with $s$ taking only the boundary and critical interior values, the minimum recorded value was given by Eq.
-77.
+78.
 That we were able to evaluate $m_{\infty}$ in Eq. 75 by restricting our search for the infimum of $L_{\infty}(\alpha,\beta,\phi;s)$ to $1\leq s\leq 30$ is due to the following.
@@
 To examine the large-$s$ regime numerically in a stable way, we re-expressed $L_{\infty}(\alpha,\beta,\phi;s)$ in terms of exponentials with negative exponent; for example,
-$\left(\frac{s}{e^{s}-1}\right)^{\beta}\left\{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}\right\}\\ =\left(\frac{s}{1-e^{-s}}\right)^{\beta}e^{(\alpha-\beta)s}\left\{(1+e^{-s})^{\alpha}-\left(1+e^{-2s}-2e^{-s}\cos(\phi)\right)^{1/2}\right\},$
+$\left(\frac{s}{e^{s}-1}\right)^{\beta}\left\{(e^{s}+1)^{\alpha}-R(s,\phi)^{\alpha}\right\}\\ =\left(\frac{s}{1-e^{-s}}\right)^{\beta}e^{(\alpha-\beta)s}\left\{(1+e^{-s})^{\alpha}-\left(1+e^{-2s}-2e^{-s}\cos(\phi)\right)^{1/2}\right\},$
 which appears in Eq.
 76 via Eqs.
-24, 32, and 59.
+25, 33, and 60.
 We then minimized the resulting large $s$ stable form of $L_{\infty}(\alpha,\beta,\phi;s)$ over $\alpha,\beta\in(0,1)$, $\phi\in[\phi_{0},\pi)$ and $s\in[30,S]$ using Maple’s global search (Optimization[Minimize]) (using high precision arithmetic), for various values of $S\in[100,10^{5}]$.
 In every run, the optimizer returned the same minimum at the lower boundary $s=30$, with the value given in Eq.
-81.
+82.
 This provides a practical numerical certification that the infimum defining $m_{\infty}$ is achieved within the finite window $1\leq s\leq 30$, bearing in mind, from Eqs.
-71 and 75, that $L_{\infty}(\alpha,\beta,\phi;s)=\mathcal{O}(s^{-1})$ as $s\to\infty$.
+72 and 76, that $L_{\infty}(\alpha,\beta,\phi;s)=\mathcal{O}(s^{-1})$ as $s\to\infty$.
 This is illustrated by our optimization computation for $s=10^{5}$ giving the value
@@
 Finally, we also used Optimization[Minimize] to independently verify the grid results that yielded Eqs.
-76 and $7$ for $s\in[0,1]$ and $s\in[1,30]$, respectively.
+77 and 78 for $s\in[0,1]$ and $s\in[1,30]$, respectively.
 The Maple source code for the evaluation of $m_{0}$ and $m_{\infty}$ is provided in [Dunster:2026:BWM].
@@
 Let $w_{n}(\alpha,\beta,x)$ be the scaled hypergeometric function defined by Eq.
-12, $A_{1,2}(\alpha,\phi)$ be given by Eqs.
-52 and 53, and $m_{0}$ and $m_{\infty}$ be the infima given by Eqs.
-76 and 77.
+10, $A_{1,2}(\alpha,\phi)$ be given by Eqs.
+53 and 54, and $m_{0}$ and $m_{\infty}$ be the infima given by Eqs.
+77 and 78.
 Then from Eq.
-28, the scaling Eq.
+29, the scaling Eq.
 56 together with Euler’s reflection identity Eq.
 8, and the decomposition Eqs.
-54 and 56 combined with the bounds Eqs.
-78 and 79, we deduce the following.
+55 and 57 combined with the bounds Eqs.
+79 and 80, we deduce the following.
 Theorem 4.1.
-For $n=1,3,5,\dots$, $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi)$
+For $n=1,3,5,\dots$, $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi]$
 (84) $w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\geq Q_{n}(\alpha,\beta,\phi)\,P_{n}(\alpha,\beta,\phi),$
@@
 and
-(85) $Q_{n}(\alpha,\beta,\phi)=\frac{\pi\,(\alpha+\beta)\sin(\pi\,\alpha)n!(\pi-\phi)^{2}}{(1-\alpha)\Gamma(\beta)\Gamma(n-\alpha)n^{1+\alpha-\beta}}.$
+(86) $Q_{n}(\alpha,\beta,\phi)=\frac{\pi\,(\alpha+\beta)\sin(\pi\alpha)\,\Gamma(1+\alpha)\,(\pi-\phi)^{2}}{(1-\alpha)\Gamma(\beta)\,n^{1+\alpha-\beta}}\geq 0.$
 Note that at the boundaries
@@
 85 with $A_{1}(\alpha,\pi)$ and $A_{2}(\alpha,\beta,\pi)$ being given by Eqs.
-62 and 65, and incidentally from this (for $n=5$) we find numerically that
+63 and 66, and incidentally from this (for $n=5$) we find numerically that
 (91) $\inf_{\alpha,\beta\in(0,1)}P_{5}(\alpha,\beta,\pi)=P_{5}(0,1,\pi)=\frac{9}{80\pi}-\frac{|m_{0}|}{5}-\frac{37\,|m_{\infty}|}{125\,e^{5}}=0.01621575275\cdots.$
+Figure 3: Graph of $P_{5}(\alpha,\beta,\phi_{0})$ for $\alpha,\beta\in[0,1]$.
 For $n=5$ the desired positivity of the infimum is given by the following.
@@
 (92) $\inf_{\begin{subarray}{c}\alpha,\beta\in(0,1)\\ \phi\in[\phi_{0},\pi)\end{subarray}}P_{5}(\alpha,\beta,\phi)=P_{5}(0,1,\phi_{0})\\ =\frac{1}{\pi(\pi-\phi_{0})^{2}}\left[\ln\left\{\csc\left(\tfrac{1}{2}\phi_{0}\right)\right\}-\frac{9\left\{\csc\left(\tfrac{1}{2}\phi_{0}\right)-1\right\}}{100}\right]\\ -\frac{|m_{0}|}{5}-\frac{37\left|m_{\infty}\right|}{125\,e^{5}}=0.001500310752\cdots.$
-Figure 3: Graph of $P_{5}(\alpha,\beta,\phi_{0})$ for $\alpha,\beta\in[0,1]$.
 Figure 4: Graph of $P_{5}(\alpha,\beta,\phi_{0})$ for $\alpha\in[0,0.1]$ and $\beta\in[0.9,1]$.
@@
 In addition, the infima $m_{0}$ and $m_{\infty}$, as given by Eqs.
-59, 72, 73, 74, and 75, also depend on $\phi_{0}$, and they too approach -$\infty$ as $\phi_{0}\to 0^{+}$.
+60, 73, 74, 75, and 76, also depend on $\phi_{0}$, and they too approach -$\infty$ as $\phi_{0}\to 0^{+}$.
 Hence from Eq.
@@
 We fixed $\phi_{0}=0.061$ and $n=5$, and used the numerical values of $m_{0}$ and $m_{\infty}$ from Eqs.
-76 and 77.
+77 and 78.
 The search sampled $\alpha$ and $\beta$ over $60^{2}$ point pairs on a grid biased toward 0 and $1$ (including boundary values).
@@
 Since the cases $\alpha,\beta\in\{0,1\}$ and $\phi=\pi$ are removable singularities, near-edge values were snapped to $\alpha,\beta=0$ or $1$, and $\phi=\pi$, with both $P_{5}(\alpha,\beta,\phi)$ and $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi$ being evaluated via the corresponding exact limiting formulas on edges and corners (see Eqs.
-62, 65, 86, 87, 88, 89, and 90).
+63, 66, 87, 88, 89, 90, and 91).
 Over the resulting candidate set in $\phi$, the smallest value recorded on the $(\alpha,\beta)$ grid yielded Eq.
-91, attained as stated at $(\alpha,\beta,\phi)=(0,1,\phi_{0})$.
+92, attained as stated at $(\alpha,\beta,\phi)=(0,1,\phi_{0})$.
 The exact value shown comes from Eq.
-87.
+88.
 Again, as an independent check, we used the Maple routine Minimize over the box $\alpha,\beta\in[0,1]$, $\phi\in[\phi_{0},\pi]$, with Digits set very high, to verify Eq.
-91.
+92.
 Finally, in [Dunster:2026:BWM] surface plots of $P_{5}(\alpha,\beta,\phi)$ for $\alpha,\beta\in[0,1]$, as well as for $(\alpha,\beta)$ near $(0,1)$, animated over $\phi_{0}\leq\phi\leq\pi$, provide further visual confirmation.
@@
 The purpose of this section is to prove the following, which in conjunction with Eqs.
-83, 91, and 95, proves that $w_{n}(\alpha,\beta,-1)-|w_{n}(\alpha,\beta,e^{i\phi})|\geq 0$ for $\phi\in[\phi_{0},\pi]$ and $n=7,9,11,\ldots$.
+84, 92, and 96, proves that $w_{n}(\alpha,\beta,-1)-|w_{n}(\alpha,\beta,e^{i\phi})|\geq 0$ for $\phi\in[\phi_{0},\pi]$ and $n=7,9,11,\ldots$.
 Theorem 5.1.
@@
 95 that $c_{2}(\alpha,\beta)$, which appears in the third term on the RHS of Eq.
-99, can take either sign, and so we split the analysis into the two cases $c_{2}(\alpha,\beta)\geq 0$ and $c_{2}(\alpha,\beta)<0$.
+100, can take either sign, and so we split the analysis into the two cases $c_{2}(\alpha,\beta)\geq 0$ and $c_{2}(\alpha,\beta)<0$.
 5.1 Case 1: $c_{2}(\alpha,\beta)\geq 0\;(\alpha\geq\beta)$
@@
 (103) $F_{1}(\alpha,\beta,\phi)=\frac{1}{\nu}\left[A_{1}(\alpha,\phi)-\frac{\max\{1-\nu,0\}}{5^{\nu}}A_{2}(\alpha,\beta,\phi)\right].$
-Since $A_{1}(\alpha,\phi)$ and $A_{2}(\alpha,\beta,\phi)$ are smooth in $(\alpha,\beta,\phi)$ and from Eqs. 60 and 63
+Since $A_{1}(\alpha,\phi)$ and $A_{2}(\alpha,\beta,\phi)$ are smooth in $(\alpha,\beta,\phi)$ and from Eqs. 61 and 64
 $A_{1}(\alpha,\phi)-A_{2}(\alpha,\beta,\phi)(1-\nu)5^{-\nu}=\mathcal{O}(\nu)\quad(\nu\to 0),$
@@
 $\nu A_{2}(\alpha,\beta,\phi)c_{2}(\alpha,\beta)\,n^{-1-\nu}\geq-\nu A_{2}(\alpha,\beta,\phi)|c_{2}(\alpha,\beta)|\,5^{-1-\nu}.$
-Therefore, in conjunction with Eqs. 99 and 100, it is evident that
+Therefore, in conjunction with Eqs. 100 and 101, it is evident that
 (104) $\partial J(\alpha,\beta,\phi;n)/\partial n\geq F_{2}(\alpha,\beta,\phi)\quad(n\geq 5,\,0\leq\alpha<\beta\leq 1),$
@@
 Thus, in summary, to establish $\partial J/\partial n\geq 0$ for $n\geq 5$ it is sufficient from Eqs.
-101 and 103 to check that $F_{1}(\alpha,\beta,\phi)\geq 0$ for $0\leq\beta\leq\alpha\leq 1$, and $F_{2}(\alpha,\beta,\phi)\geq 0$ for $0\leq\alpha<\beta\leq 1$.
+102 and 104 to check that $F_{1}(\alpha,\beta,\phi)\geq 0$ for $0\leq\beta\leq\alpha\leq 1$, and $F_{2}(\alpha,\beta,\phi)\geq 0$ for $0\leq\alpha<\beta\leq 1$.
 The following then completes the proof of Theorem 5.1.
@@
 attained at $(\alpha,\beta,\phi)=(\alpha_{2},1,\phi_{0})$, where $\alpha_{2}=0.8583872779\cdots$. In particular, both infima are strictly positive.
-The computation of these values was performed in a similar manner to the method used for Proposition 4.2, that is, sampling over a fine $\alpha,\beta$ grid, and at $\phi$ boundary points ($\phi_{0}$ and $\pi$) and at any interior critical points.
+The computation of these values was performed in a similar manner to the method used for Proposition 4.2, that is, sampling over a fine $\alpha,\beta$ grid, and $\phi$ only at boundary points ($\phi_{0}$ and $\pi$) and at any interior critical points.
 The infima were then independently confirmed using the Maple Optimization[Minimize] package.
@@
 (110) $F_{2}(\alpha,1,\phi)=\frac{1}{1+\alpha}\left[A_{1}(\alpha,\phi)-\frac{(1-\alpha)\,C_{2}(1,\phi)}{5^{1+\alpha}\,\pi}\left\{\alpha+\frac{1}{10}(1+\alpha)^{2}(1-\alpha)\right\}\right],$
-where from Eqs. 23 and 31
+where from Eqs. 24 and 32
 (111) $C_{2}(1,\phi)=\frac{\csc\left(\tfrac{1}{2}\phi\right)-1}{2(\pi-\phi)^{2}}.$
@@
 (113) $F_{2}(\alpha,\beta,\pi)=\frac{1}{\nu}\left[A_{1}(\alpha,\pi)-\frac{A_{2}(\alpha,\beta,\pi)}{5^{\nu}}\left(|1-\nu|+\frac{\nu\left|c_{2}(\alpha,\beta)\right|}{5}\right)\right],$
-where $A_{1}(\alpha,\pi)$, $A_{2}(\alpha,\beta,\pi)$ and $c_{2}(\alpha,\beta)$ are given by Eqs. 62, 65, and 94.
-Incidentally, in comparison to Eqs. 105 and 106, from Eqs. 111 and 112 we find
+where $A_{1}(\alpha,\pi)$, $A_{2}(\alpha,\beta,\pi)$ and $c_{2}(\alpha,\beta)$ are given by Eqs. 63, 66, and 95.
+Incidentally, in comparison to Eqs. 106 and 107, from Eqs. 112 and 113 we find
 $\inf_{\alpha\geq\beta}F_{1}(\alpha,\beta,\pi)=F_{1}(1,1,\pi)=\frac{1}{8\pi}>\inf_{\alpha<\beta}F_{2}(\alpha,\beta,\pi)=0.03287693288\cdots,$
@@
 28].
-We finish by noting the following simple result, which confirms without using Kummer functions that Eq. 5 holds at $\theta=\pi$ ($\phi=0$) provided $\beta\geq\alpha$.
+We finish by noting the following simple result, which confirms without using Kummer functions that Eq. 5 holds ($n$ even or odd) at $\theta=\pi$ ($\phi=0$), provided $\beta\geq\alpha$.
 Lemma 6.1.
-Let $0<\alpha\leq\beta<1$ and let $1\leq n\in\mathbb{N}$ be odd. Then
+Let $0<\alpha\leq\beta<1$ and let $1\leq n\in\mathbb{N}$. Then
 (121) $A_{n}(\alpha,\beta,1)>\bigl|A_{n}(\alpha,\beta,-1)\bigr|.$
@@
 (124) $\binom{\alpha}{k}=\frac{\Gamma(1+\alpha)}{\Gamma(1-k+\alpha)\,\Gamma(k+1)}\\ =(-1)^{k+1}\frac{\Gamma(1+\alpha)\,\Gamma(k-\alpha)\sin(\pi\alpha)}{\pi k!}\quad(k=1,2,3,\ldots n).$
-Thus for $\alpha\in(0,1)$ the $k=0$ and the $k$-odd coefficients in Eq.
-122 are positive.
-Therefore, using Eq.
-122 with $\omega=\pm 1$ and the fact that $n$ is assumed to be odd, we obtain for $0<\alpha\leq\beta<1$
+Thus for $\alpha\in(0,1)$ the $k=0$ and the $k$-odd coefficients in Eq. 123 are positive. Therefore, using Eq. 123 with $\omega=\pm 1$, we obtain for $0<\alpha\leq\beta<1$
 $A_{n}(\alpha,\beta,1)-A_{n}(\alpha,\beta,-1)=A_{n}(\alpha,\beta,1)-\left|A_{n}(\alpha,\beta,-1)\right|\\ =\sum_{k=0}^{n}\binom{\alpha}{k}\,\frac{\Gamma(n-k+\beta)}{\Gamma(\beta)\,\Gamma(n-k+1)}\left\{1-(-1)^{k}\right\}=2\sum_{\begin{subarray}{c}1\leq k\leq n\\ k\ \mathrm{odd}\end{subarray}}\!\!\binom{\alpha}{k}\,\frac{\Gamma(n-k+\beta)}{\Gamma(\beta)\,\Gamma(n-k+1)}>0.$